### v1

# Assist Research

Abstract
Numerous studies have assessed the proficiency of AI systems, particularly large language models (LLMs), in facilitating everyday tasks such as email writing, question answering, and creative content generation. However, researchers face unique challenges and opportunities in leveraging LLMs for their own work, such as brainstorming research ideas, designing experiments, and writing or reviewing papers. In this study, we introduce AAAR-1.0, a benchmark dataset designed to evaluate LLM performance in four fundamental, expertise-intensive research tasks: (i) EquationInference, assessing the correctness of equations based on the contextual information in paper submissions; (ii) ExperimentDesign, designing experiments to validate research ideas and solutions; (iii) PaperWeakness, identifying weaknesses in paper submissions; and (iv) ReviewCritique, identifying each segment in human reviews is deficient or not. AAAR-1.0 differs from prior benchmarks in two key ways: first, it is explicitly research-oriented, with tasks requiring deep domain expertise; second, it is researcher-oriented, mirroring the primary activities that researchers engage in on a daily basis. An evaluation of both open-source and closed-source LLMs reveals their potential as well as limitations in conducting sophisticated research tasks. We will release the AAAR-1.0 and keep iterating it to new versions. Project Webpage:   https://renzelou.github.io/AAAR-1.0/
Figure 1: The input-output illustration of four tasks in the proposed AAAR-1.0 benchmark.
1 Introduction
Although AI has brought transformative changes to various aspects of life, its impact on researchers unfolds in a nuanced manner. On the one hand, AI assists in various research disciplines, such as Social Science (Neuman et al., 2023), Finance (Gu et al., 2024), Medicine (Rakhimov et al., 2022), GeoScience (Praskievicz, 2018), etc., significantly expediting academic processes. However, many of these applications are superficial, often limited to data-driven clustering or classification. On the flip side, the AI era poses challenges for researchers. Despite its ability to streamline some activities, researchers still face demanding, cognitively intensive tasks such as staying current through extensive paper reading, rapidly generating ideas in response to fast-paced advancements, conducting rigorous experiments to substantiate claims, and managing an increasing volume of peer reviews. Then a question looms: How effectively can AI assist researchers in tasks that are domain-specific, expertise-demanding, and reasoning-intensive?
Existing works proved the promising potential for using LLMs in assisting AI research. Si et al. (2024) conducted a large-scale human study and found that LLMs can generate creative research ideas. Lu et al. (2024) proposed an autonomous agent to handle complicated research workflow and write a whole research paper. However, most of these works focus on addressing highly subjective problems that require a high degree of expertise, making evaluation laborious and hard to reproduce. This underscores the need for a comprehensive benchmark that rigorously assesses LLMs’ capabilities in expertise-intensive research activities.
To this end, in this work, we introduce AAAR-1.0, a novel benchmark that aims to comprehensively assess the LLMs’ capacity on expert-level research tasks. As illustrated in Figure 1, AAAR-1.0 decomposes four distinct expert-level AI research tasks from the researcher’s daily activities, including i) EquationInference, investigating whether the LLMs can infer the equation correctness based on the paper context; ii) ExperimentDesign, validating LLMs’ ability on designing reliable experiments for a research idea; iii) PaperWeakness, testing the quality of weaknesses discovered by LLMs from paper drafts; and iv) ReviewCritique, investigating whether LLMs can identify and explain the deficient/unreliable human-written paper reviews. To ensure data quality, senior AI researchers with extensive domain expertise perform data annotation for AAAR-1.0, followed by rigorous multi-round data examination and filtering. All four tasks require models to possess strong domain knowledge covering various cutting-edge research findings, as well as expert-level research experience, to the extent that even humans need substantial research accumulation to tackle the tasks we designed. Crucially, tasks here are singular, stand-alone challenges (with clear input and output expectations) rather than a complicated task chain (Li et al., 2024; Lu et al., 2024), providing a more transparent assessment of the model’s intermediate output.
Benefiting from the proposed automatic metrics, we conduct extensive experiments across numerous mainstream LLMs, where we find that:
  • With a random guess baseline of 25%, the performance of most LLMs on EqInfer hovers just slightly above chance, with the top models reaching around 60%. This highlights the difficulty of the task, despite its reliance only on local context reasoning.
  • In ExpDesign, LLM-designed experiments are innovative and more diverse than those by humans; however, many are trivial, lack feasibility, and stray from the original research objectives.
  • In PaperWeakness, LLM-identified weaknesses often lack depth and specificity, making them broadly applicable and less useful for providing feedback on paper drafts.
  • In ReviewCritique, LLMs struggle to effectively identify deficient human reviews, indicating limited usefulness in assisting meta-reviewers in evaluating the quality of individual paper reviews.
2 Related Work
LLMs for AI Research.
With the rapid evolution of pertaining techniques, LLMs are found to be useful in assisting various research disciplines (Yu et al., 2024; Labrak et al., 2024), particularly in AI research, such as generating novel research ideas (Kumar et al., 2024), reviewing research draft (Gao et al., 2024; Du et al., 2024; Liang et al., 2024), and writing scientific papers (Chamoun et al., 2024; Lu et al., 2024). For example, Si et al. (2024) conducted a large-scale human investigation on LLM-generated research ideas and found that LLMs can generate novel ideas compared with humans while lacking feasibility. Du et al. (2024) found that while LLMs are effective at summarizing papers, they tend to overly trust the authors’ claimed strengths and struggle to identify weaknesses specific to the paper. Furthermore, some works try to employ LLMs to solve more complicated research tasks that are composed of multiple steps (Li et al., 2024; 2023; Tang et al., 2023). Notably, Lu et al. (2024) proposed AI-Scientist, an autonomous agent framework that can handle a series of challenging research tasks consecutively, including generating research ideas, coming up with the corresponding experiments along with the implementations, and then writing the final research paper — exactly how human conduct a whole research pipeline. However, there is still a lack of systematic evaluations and quantitative analyses on the LLMs’ (intermediate) output of each single-step research task. Our work focuses on building a benchmark that has individual research steps with clear input-output expectations, thus making it suitable for comprehensive LLMs evaluation.
Benchmarks for AI Research Tasks.
Existing “LLM assists research” benchmarks mainly focus on the implementation and execution part of the research pipeline (Lu et al., 2024; Chen et al., 2024b; Li et al., 2024). For instance, Huang et al. (2024) proposed MLAgentBench to test the LLMs’ capacity for writing project code and training the ML models, where the evaluation metric is the test performance of the models trained by LLMs. However, real-world AI research activities are diverse and some of them are hard to assess for quality, such as generating research ideas, which requires intensive manual assessment (Si et al., 2024; Liang et al., 2024), or LLM-based estimation (Lu et al., 2024). Our work centers on tasks that emphasize a comprehensive mastery of the scientific research field and core elements of a researcher’s daily workload, and we try to build curated task-specific metrics for every single task for a more efficient and accurate LLMs appraisal.
Figure 2: Data construction workflows of the three tasks in Refer to caption AAAR-1.0.
3 AAAR-1.0
Figure 2 provides an overview of the construction process of AAAR-1.0. In the following sections, we elaborate on the data collection details of the aforementioned four tasks, including § 3.1 EquationInference (EqInfer), § 3.2 ExperimentDesign (ExpDesign), § 3.3 PaperWeakness (Weakness), and § 3.4 ReviewCritique.
3.1 EquationInference
Crafting a correct scientific equation in paper writing or validating an equation in paper reviewing is challenging, as it requires a thorough understanding of an algorithm or the intricate relationships among numerous variables. Directly prompting LLMs to generate equations proves overly demanding. Therefore, this work formulates EqInfer (Figure 1) as as a multiple-choice equation inference task based on contextual cues, curated as follows.
① Data crawling and cleaning.
For the data source, we adopt the pre-compilation LaTeX code for two reasons: i) existing PDF parsing tools, such as PyMuPDF and PaperMage (Lo et al., 2023), can introduce considerable noise to the parsed equation text; ii) considering most of exiting LLMs are capable with processing LaTeX code, using LaTeX source instead of parsed text can be more accurate and provide LLMs with richer information. Meanwhile, to avoid using any low-quality human-written equations, we only crawl those peer-reviewed papers accepted by top-tier conferences. Accordingly, we first obtain the accepted paper list from ACL Anthology, from year 2019 to 2023. Next, we search each paper on arXiv to crawl its LaTeX source (if it exists). Finally, we get a total of 1,762 papers’ source LaTeX packages.
We then clean the LaTeX sources by deleting all the comments and combining multiple cross-referred .tex files into a main file. Afterwards, we use regex to randomly extract (at most) 3 equations’ code snippets per paper, finally resulting in 3,877 human-written equations are extracted.
② LLM-based equation synthesis.
As we formulate this task as classification, for each human-written positive equation, we have to craft at least three counterpart negative equations. To this end, we prompt GPT-4 to synthesize more equations based on the paper context. For each positive equation, we repeat this prompting (with a high decoding temperature) until three different negative equations are synthesized.
③ LLM-based filtering.
However, the LLM-synthetic equations can sometimes be context-unaligned, i.e., some synthesized equations contain notations that are never defined in the paper context, which is a superficial shortcut for the classification tasks. To improve the data quality, we prompt GPT-4 to identify those context-unaligned negative equations. We then discard those instances where all three negative equations are identified as contextually unaligned. This filtering leads to a final of 1,449 classification instances (62.3% instances are filtered).
④ Expert-based examination.
Furthermore, it’s also possible that synthesized negative equations are actually correct (i.e., false negative options) — even if the negative and positive equations are written differently, the final compiled results might be the same. To filter out the false negative equations and to have a final check on the classification instances, we then employ human experts to conduct a further data review.
We asked 5 senior Ph.D. students who are experienced in AI research to manually check all the instances. For each classification instance, we ask those human experts to consider the following criteria: i) Are all four equations (both positive and negative) grammatically correct? ii) After compilation, is there only one correct answer? We ask every human expert to use external LaTeX compilation tools (e.g., TeXlive), and identify the instances that cannot meet the criteria. Each instance is examined by at least two experts, and we only keep instances that all experts decide to keep. After this strict examination, a total of 1,049 instances are eventually kept (27.6% instances are filtered)
Final data.
We finally shuffle the four equations for each classification instance and randomly assign letters (A, B, C, and D) to the equations. We show the data statistics of the final EqInfer in Table 11 and the sample data cases in Figure 7.
3.2 ExperimentDesign
Given a research topic, such as a novel ML algorithm, a qualified researcher can design a solid experiment plan for it, and clarify underlying motivation to ensure the reliability of the designed experiment. Unlike the concurrent works that focus on the experiment implementation (Lu et al., 2024; Huang et al., 2024), we emphasize the importance of assessing the high-level experiment design of LLMs before the subsequent implementation to avoid any expensive execution iteration. Therefore, as shown in Figure 1, we formulate ExpDesign as a text-generation task that takes pre-experiment paper context as input, and then generates the experiment and explanation list.
① Data crawling.
As for the data source, we first collect $\geq$ 10k papers’ data from arXiv, including LaTeX sources and PDFs, which cover broad AI categories, including cs.AI, cs.CL, and cs.CV, from year 2018 to 2023. Similarly, to ensure the source data quality, we only use papers that have appeared at well-known conferences.
② Domain-expert annotation.
Making a reliable and executable experiment plan requires solid foundation knowledge of a specific research area. Consequently, we set a high standard for choosing annotators: i) be a senior Ph.D. student with at least one peer-reviewed publication in leading AI venues; ii) have more than 4 years of AI research experience; iii) frequently serve as conference reviewers. Finally, we invite a total of 10 qualified experts to participate in our data collection procedure. Given the 10k crawled papers, we first ask every annotator to bid on the papers that they are interested in. After bidding, each of them is assigned 10 papers, i.e., a total of 100 papers to be annotated. During annotation, we post each paper PDF on online Google Drive and ask the annotator to first carefully read the whole paper. Then, we ask them to identify and locate the key experiments in each paper (i.e., highlighting the relevant paragraphs of each experiment). We don’t consider some trivial experiments, such as those supplemental analyses in the appendix section. For each identified experiment, the annotator has to concisely answer two questions: i) What did this experiment do? ii) Why did the paper authors conduct this experiment? In other words, we ask the annotator to summarize all the key experiments in this paper and explain the underlying motivations based on their rich domain experience.
③ Multi-round peer discussion.
Intuitively, different experts might have different opinions on the same research topic. Particularly, when explaining the underlying motivation of an experiment, adopting only a single expert’s opinion might introduce bias to our annotation. Hence, we conduct a further multi-round peer discussion. For each online paper PDF, where all the key experiments are identified, summarized, and explained, we ask a different expert (reviewer) to review the annotation by considering the following three criteria: i) Are the identified experiments all the key experiments? ii) Does each experiment summarization covers all key information? iii) Does each explanation sound reasonable and reliable? Each reviewer has to leave comments to the online PDF regarding the above criteria, and then the annotator has to respond to each comment — either accept the suggestion and revise the previous annotation, or provide a “rebuttal” to the reviewer to uphold the annotation. This discussion iterates until both opinions align with each other. Eventually, for each paper, we collect two lists: i) the experiment list, summarizing each experiment step of the paper; ii) the explanation list, the underlying motivations that are one-one corresponding to the experiment.
Final data.
After annotation, we use the pre-experiment context of each paper (according to the first-experiment location identified by the annotator) as the input. Furthermore, we use GPT-4 to delete any sentence that potentially leaks the experiment from the input. [^1] About 9.8% sentences are deleted. Similar to the EqInfer, we utilize the source LaTeX as the input text to avoid PDF paring noise. As for the image input, we collect those figures within each paper’s source LaTeX package and only keep figures that are used in the pre-experiment context. Overall, a total of 100 instances are collected. As shown in Figure 1, the input of each instance is the pre-experiment context (including the figures), and the ground-truth output is the expert-annotated experiment plan and the explanations. Table 12 shows data statistics and Figure 8 illustrates the sample case in ExpDesign.
3.3 PaperWeakness
Another critical research task is paper review. Previous works have demonstrated the usefulness of the LLM-based review feedback (Gao et al., 2024; Jin et al., 2024; Lu et al., 2024). However, as indicated by Du et al. (2024); Liang et al. (2024), LLMs only excel at summarizing the research strengths while falling significantly short on weakness criticism. Hence, we build Weakness for particularly investigating the LLM-generated weaknesses.
① Data crawling.
We first crawl a total of 3,779 anonymous submissions of ICLR 2023 from OpenReview, [^2] We adopt ICLR because it releases full submissions, while some other conferences only release accepted papers. including PDF and other meta information (e.g., scores, decisions, and tracks). As the ICLR 2023 has 13 distinct tracks while the paper distribution across different tracks is highly biased, we then uniformly sample papers from different research tracks to improve the domain diversity. Meanwhile, during sampling, we also keep the accept/reject papers distributed equally to avoid data bias. In a word, we finally collect a total of 1,000 papers (500 accepted; 500 rejected), uniformly covering all 13 tracks. Please refer to Figure 5 for the track and score distribution of the 1,000 papers.
② Extraction of human-written weaknesses.
Since the raw comments crawled from ICLR 2023 are mixed with both strengths and weaknesses, we further employ GPT-4 to extract all the weaknesses from each reviewer’s comments and compose multiple weaknesses into a list. Notably, we force GPT-4 to keep the original text of the reviewer, i.e., all weaknesses in our dataset are those original sentences written by the reviewer without any modifications. [^3] We manually checked GPT-4’s extraction results of 200 cases — GPT-4 only missed $\leq$1of reviewer-written weaknesses and maintained almost all the original text. What’s more, sometimes one reviewer might repeatedly mention the same weakness throughout the comment. In this case, we simply keep all the repeated weaknesses because, if one weakness is repeatedly mentioned by the reviewer, it’s intuitively an important weakness that the reviewer wants to emphasise; accordingly, keeping the repeat items can penalize LLMs more on missing this weakness.
For each paper, we can finally get multiple weakness lists (one weakness list per reviewer, one paper can have multiple reviewers). We further delete a few papers without any weaknesses found in the raw comments, resulting in a total of 993 instances, i.e., 993 {paper, weakness lists} pairs.
③ Input data processing.
As we mentioned before, we crawl papers from OpenReview instead of arXiv because the under-review paper draft is required for this task. However, not every paper from OpenReview can be found on arXiv, i.e., the source LaTeX code and figures of most under-review papers are unavailable. Therefore, we utilize VILA (Lin et al., 2023) to parse text data out from the PDF; we also employ PDFFigures-2.0 (Clark & Divvala, 2016) to extract all the figures and tables (in image) from the paper, as Vila is not good at processing the table data.
Final data.
Our final data is composed of 993 instances, each input is paper text along with figure/table images, and each output is peer reviewers’ weakness lists. Table 13 shows data statistics; Figure 9 presents an example of the data instances. We show the data diversity (score and track distribution) in Figure 5.
3.4 ReviewCritique
In addition to identifying weaknesses in paper drafts, a more challenging research task that requires more senior research experience is conducting meta-reviewing. Given a paper submission, along with individual reviews and author rebuttals, meta-reviewing is not to summarize individual reviews. Instead, a meta-reviewer must go through all the information and make a final recommendation. This requires the meta-reviewer to identify deficient/unreliable review segments (e.g., if a viewpoint is too subjective or contains factual errors) in each individual review and make a decision based on the non-deficient ones. This task demands years of experience in the relevant domain; even for human experts, only senior researchers are typically qualified for meta-reviewing. Therefore, as illustrated in Figure 1, we also investigate how LLMs assist meta-reviewers, specifically in identifying deficient review points.
We reuse the ReviewCritique dataset from our recent work (Du et al., 2024), where we crawled papers’ initial submissions along with their reviews from OpenReview and employed more than 40 AI research experts to label each review segment (i.e., deficient or not), with detailed human explanations. In total, there were 100 papers with 380 human reviews. Each review was divided into sentence-level segments, resulting in 11,376 review segments (viewpoints).
4 Evaluation Criteria
For EqInfer, we adopt accuracy as the classification criterion. For ExpDesign and Weakness, since both tasks have natural language outputs, semantic-based metrics are necessary. Hence, in addition to the conventional ROUGE (Lin, 2004), we also develop several novel similarity-based metrics for each specific task.
When evaluating the experiment plan list of ExpDesign, we hope the LLMs can mention as many similar experiment steps as the expert’s plan. Nevertheless, we also don’t expect LLMs to generate too many irrelevant or redundant steps in the plan. This intuition covers both the “recall” and “precision” aspects. Therefore, we develop semantic similarity-based F1 score, denoted as S-F1, which is the harmonic mean of S-Precision and S-Recall:
S-Precision $=\frac{1}{m}\sum_{i=1}^{m}\max_{j}\textrm{sim}(p_{i},g_{j})$ (1)
S-Recall $=\frac{1}{n}\sum_{j=1}^{n}\max_{i}\textrm{sim}(g_{j},p_{i})$ (2)
where the $p$ and $g$ represent the LLM’s prediction plan and the ground-truth plan, respectively. The $m$ and $n$ are the list length of $p$ and $g$ (e.g., $m$ experiment steps in $p$). We use SentenceBERT (Reimers, 2019) to measure the semantic similarity between the $p_{i}$ step and the $g_{j}$ step.
Meanwhile, S-F1 omits the item order difference of two lists, but when giving same-length lists (items have one-one correspondence), we can utilize the following similarity-based matching score:
[…]
Unlike ExpDesign, the output of Weakness is multiple reviewers’ weakness lists, which means we have to measure LLM’s single prediction list with a “nested” list. Hence, we rewrite S-Precision ,  S-Recall to SN-Precision , SN-Recall:
SN-Precision $=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{r}\sum_{k=1}^{r}\max_{j}\textrm{sim}(p_{i},g^{k}_{j})\right)$ (4)
SN-Recall $=\frac{1}{r}\sum_{k=1}^{r}\left(\frac{1}{n_{k}}\sum_{j=1}^{n_{k}}\max_{i}\textrm{sim}(g^{k}_{j},p_{i})\right)$ (5)
where $r$ is the number of reviewers of the given paper, $n_{k}$ means the length of $k$-th reviewer’s weakness list, and $g^{k}_{j}$ indicates the $j$-th item in $k$-th reviewer’s weakness list.
Additionally, in the real world, we would think a review weakness is reliable if it is specific to a paper. Meanwhile, we also hope the review is informative, i.e., no excessive similar weaknesses in one review. Therefore, we adopt the following ITF-IDF metric proposed by Du et al. (2024), which is inspired by the classic TF-IDF:
ITF-IDF $=\frac{1}{w}\sum_{j=1}^{w}\left(\frac{1}{m_{j}}\sum_{i=1}^{m_{j}}\log\left(\frac{m_{j}}{O_{i}^{j}}\right)\times\log\left(\frac{w}{R_{i}^{j}}\right)\right)$ (6)
$O_{i}^{j}$ $=\sum_{k=1}^{m_{j}}\textrm{sim}(p_{i}^{j},p_{k}^{j})$ (7)
$R_{i}^{j}$ $=\sum_{l=1}^{w}\max_{s}\textrm{sim}(p_{i}^{j},p_{s}^{l})$ (8)
where the $w$ is the total number of papers in the dataset, $p^{j}$ is $j$-th paper’s prediction weakness list, $p^{j}_{i}$ is the $i$-th weakness in $p^{j}$. Moreover, ${O_{i}^{j}}$ calculates the intra-paper occurrence frequency of $p^{j}_{i}$; ${R_{i}^{j}}$ is the “soft” number of papers that also contain the $p^{j}_{i}$, which is computed by summing the maximum similarity scores between $p^{j}_{i}$ and other paper’s weaknesses. In a word, ${O_{i}^{j}}$ measures informativeness, and ${R_{i}^{j}}$ measures specificity. The complete ITF-IDF consider both aspects and reflects the overall weakness diversity.
For ReviewCritique, we use F1 score as the classification metric; while for the deficiency explanation, we use ROUGE (Lin, 2004) and BERTScore (Zhang et al., 2020) to reflect how well the model-generated explanation aligns with the expert’s annotation.
5 Experiments and Analyses
In this section, we conduct extensive experiments on AAAR-1.0, across various mainstream LLMs, to quantify the current LLMs’ capacity to tackle high-level research tasks. Specifically, § 5.1 for EqInfer, § 5.2 for ExpDesign, § 5.3 for Weakness, and § 5.4 for ReviewCritique. Please refer to Appendix B.2 for running details of the LLMs.
5.1 EquationInference
Table 1: Various LLMs’ performances on the 1,049 instances of EqInfer task.
Methods Accuracy (%)
Random Guess 25.00
Open-source LLMs
Gemma 2-27B (Gemma Team, 2024) 3.24
Falcon-40B (Almazrouei et al., 2023) 4.39
OLMo-7B (Groeneveld et al., 2024) 19.00
Mistral-7B (Jiang et al., 2023) 22.21
Qwen 2.5-72B (Qwen Team, 2024) 35.93
Mixtral-8x22B-MoE (Jiang et al., 2024) 37.08
Llama 3.1-70B (MetaAI, 2024) 38.13
Closed-source LLMs
Gemini 1.5 Pro (Anil et al., 2023) 34.31
GPT-4o (OpenAI, 2024a) 43.18
GPT-4 (OpenAI et al., 2023) 49.85
o1-preview (OpenAI, 2024b) 59.49
Claude 3.5 sonnet (Anthropic, 2024a) 61.10
Figure 3: The input context length scaling trend on the EqInfer task.
Settings.
As different LLMs have distinct context windows, to ensure a fair comparison, we fix the maximum input length for all models. According to the data statistics of Table 11, we empirically use 1,000 words for both contexts before and after equations, i.e., 2,000 surrounded words.
Main results.
Table 1 shows the main results. Firstly, the open-source LLMs, especially the Falcon and Gemma, perform unexpectedly disappointing (even worse than random guesses). These screwed scores are mainly due to the poor long-context instruction following ability, where we find some open-source LLMs are confused with the massive input and often copy the LaTeX code from the input. In contrast, closed-source LLMs generally achieve superior accuracy, probably owing to the richer scientific knowledge from the larger model parameters. However, considering the conventional multi-choice QA formulation of EqInfer, the recently-released GPT-4o solely gets 43.18, implying the unique challenge of EqInfer compared with other scientific QA benchmarks (Song et al., 2023). Notably, with the help of internal CoT, o1 gains stronger performances than GPT-4/GPT-4o, indicating the potential benefits of adopting reasoning for this task.
$\mathcal{Q}$: Do more contexts boost performance?
Table 1 unifies the input context lengths to 1,000 tokens for various LLMs. To answer this question, we experiment with long-context LLMs to investigate the impact of the input context lengths. Particularly, we scale the input length (per side) from 100 to 1,500 words. As shown in Figure 3, for the open-source LLMs (Llama and Qwen), after 300 words length, increasing the input context doesn’t help the performance and even significantly drops Qwen’s scores. While for the closed-source GPT-4-Turbo and GPT-4o, scaling up input length gradually boosts the performances at the first 1,000 words, but stabilizes afterwards. This is in line with human intuition, i.e., surrounding context is required for the equation inference, as the adjacent context usually provides important information, such as the target algorithm description or the notation definition. However, after exceeding a specific threshold, more context information is not beneficial anymore and even confuses those LLMs with poor long-context handling capacity (Wang et al., 2024; Liu et al., 2024).
5.2 ExperimentDesign
Table 2: Various LLMs’ performances on the 100 instances of ExpDesign. The explanation generation is based on the oracle experiments to prevent error propagation. “Copy Input” is a random baseline: for experiment design, randomly select 5 sentences from the input paper; for experiment explanation, directly copy each experiment idea.
Experiment Design Experiment Explanation
Methods S-F1 S-Precision S-Recall S-Match ROUGE-L ROUGE-1
Copy Input 21.13 17.94 26.76 40.32 22.06 25.28
Open-source LLMs
OLMo-7B (Groeneveld et al., 2024) 33.94 37.25 31.79 45.78 26.30 30.38
Falcon-40B (Almazrouei et al., 2023) 17.87 21.78 15.35 17.03 12.10 12.72
Gemma 2-27B (Gemma Team, 2024) 34.33 39.71 30.51 42.77 26.20 29.63
Mistral-7B (Jiang et al., 2023) 37.62 43.09 34.19 50.18 30.20 34.69
Mixtral-8x22B-MoE (Jiang et al., 2024) 42.21 50.13 36.82 49.07 29.96 34.53
Llama 3.1-70B (MetaAI, 2024) 40.57 48.43 35.43 50.05 29.33 34.11
Qwen 2.5-72B (Qwen Team, 2024) 43.24 51.73 37.55 51.12 29.46 34.68
Closed-source LLMs
Gemini 1.5 Pro (Anil et al., 2023) 51.87 50.77 53.37 52.87 28.52 33.80
Claude 3.5 sonnet (Anthropic, 2024a) 48.74 46.49 51.53 53.03 18.75 26.15
GPT-4 (OpenAI et al., 2023) 43.89 42.34 45.82 55.03 22.82 30.01
GPT-4o (OpenAI, 2024a) 53.00 51.24 55.12 54.79 27.54 34.31
o1-preview (OpenAI, 2024b) 46.67 45.04 48.70 58.55 29.11 36.70
Settings.
Similarly, we unify the input context length of different LLMs to ensure a fair comparison. According to Table 12, we set 2,000 and 3,000 input words for open- and closed-source LLMs, respectively. Meanwhile, as experiment explanation is the subsequent task of experiment design, using model-generated experiments can propagate errors in explanation, leading to inferior results for most LLMs. To this end, we provide LLMs with the oracle experiments when generating explanations.
Main results.
Table 2 shows the main results. For the experiment design, the closed-source LLMs generally outperform open-source LLMs, and both closed-/open-source LLMs are superior to the “Copy Input” baseline (except the Falcon). Despite the higher S-Precision, the open-source LLMs are seriously deficient in S-Recall compared with closed-source LLMs ($\sim$10$\downarrow$). We find that closed-source LLMs are more creative in experiment design and tend to generate more experiment ideas than open-source LLMs (though most of the experiment ideas are trivial), leading to excellent S-Recall. As for the experiment explanation, the S-Match scores of closed-source LLMs still surpass the open-source LLMs, while the score difference is not significant. Furthermore, we find the negative correlation between S-Match and the ROUGE, where the ROUGE scores of closed-source LLMs are broadly inferior. We find that the open-source LLMs often try to copy the terms or phrases from the given experiment, or even simply paraphrase the experiment instead of explaining, which results in a high superficial overlap with the ground-truth explanation. This observation highlights the importance of adopting the proposed S-Match to avoid evaluation bias of traditional generation metrics.
$\mathcal{Q}_{1}$: Can self-contained experiment design enhance the experiment explanation?
Table 3: The impact on S-Match scores of maintaining the experiment’s self-containment for ExpDesign.
Models One-by-One Whole-List
Llama 3.1-70B 50.05 49.36 ($\downarrow$ 0.7)
Qwen 2.5-72B 51.12 48.56 ($\downarrow$ 2.6)
Gemini 1.5 Pro 52.87 57.48 ($\uparrow$ 4.6)
Claude 3.5 sonnet 53.03 59.11 ($\uparrow$ 6.1)
GPT-4 55.03 56.95 ($\uparrow$ 1.9)
GPT-4o 54.79 58.54 ($\uparrow$ 3.8)
o1-preview 58.55 61.58 ($\uparrow$ 3.0)
When generating the explanation in Table 2, we provide LLMs with each individual experiment and let them explain one by one, because we find that, when providing the whole experiment list, those open-source models only explain partial experiments because of their poor instruction-following capacity. However, there are intuitively some semantic or logical relations between different experiments, e.g., some experiments are prerequisites to others. Therefore, this one-by-one prompting might break the self-containment of an experiment plan. Consequently, we test with the “whole-list” prompting, where the LLMs are given the complete experiment list and are asked to explain all experiment steps together.
As shown in Table 3, unlike the open-source LLMs, the explanation performances of those closed-source LLMs are generally improved after adopting whole-list prompting. According to further manual checking, after maintaining the self-containment of the experiments, the LLMs can refer to other experiments and better grasp the underlying motivation of the current experiment.
$\mathcal{Q}_{2}$: Do human evaluation results align with automatic metrics for explanation?
Table 4: The human evaluation results on LLMs’ output explanations of ExpDesign. “Acc. ratio” means how many model outputs are accepted by the annotator.
Models Acc. ratio
Llama 3.1-70B 22.93
Gemini 1.5 Pro 55.07
Claude 3.5 sonnet 61.46
GPT-4o 69.72
o1-preview 76.14
As the explanation can be open-ended, in this paragraph, we provide the human evaluation results on different LLMs’ experiment explanation outputs. In detail, we randomly select 20 out of 100 papers and ask 5 annotators to read the experiments along with each model’s explanations; we then let the annotator decide whether each model’s explanation is acceptable (see Appendix C.1 for more details). Table 4 illustrates the results, where the score variance is higher than Table 2. However, the performance ranking of both tables is perfectly correlated with each other (Spearman’s rank correlation coefficient $=$ 1), demonstrating the effectiveness of S-Match.
$\mathcal{Q}_{3}$: Do more contexts boost performance?
Figure 4: The input context length scaling trend of different LLMs on the ExpDesign task.
We also investigate the impact of input context length for ExpDesign. As shown in Figure 4, we scale up the input pre-experiment context length from 0.1k to 10k tokens (10k tokens is the maximum paper context length in the dataset). For the experiment planning, more input context does improve the performance of different LLMs, while this benefit stops after exceeding 5k tokens, which is similar to EqInfer’s scaling results — after the necessary information has been covered, scaling more up doesn’t boost the performance. Meanwhile, the results of the experiment explanation demonstrate that explaining motivations almost doesn’t require any paper context, i.e., the LLMs solely rely on the given experiments. However, we do not expect this because we hope LLMs can explain the motivation based on a thorough understanding of the paper, just like how human experts do. Hence, there is still a considerable gap between the LLMs and humans in terms of grasping research motivations.
$\mathcal{Q}_{4}$: Does multi-modal input boost performance?
Table 5: The figure inputs ablation of ExpDesign. For the maximum text input length, same as the setting in Table 2, we use 2,000 and 3,000 words for open- and closed-source models, respectively. For the closed-source GPT-4o and GPT-4, as they have long context window sizes, we use all the figures of each paper. While for InternVL2, we randomly select two figures per input paper.
Experiment Design Experiment Explanation
Models S-F1 S-Precision S-Recall S-Match ROUGE-L ROUGE-1
GPT-4o 53.00 51.24 55.12 58.54 29.25 35.50
w/ figures 50.11 48.94 51.59 58.53 27.87 34.30
GPT-4 43.89 42.34 45.82 56.95 25.98 33.37
w/ figures 43.54 42.56 44.85 55.03 22.82 30.01
InternVL2-26B 40.52 48.95 35.20 50.03 29.13 34.26
w/ figures 38.83 46.91 33.70 50.29 29.29 34.06
Intuitively, besides the text, when designing experiments for a given research topic, the figures can provide rich supplementary information, such as an algorithm illustration that can help better understand this research topic and underlying motivations. Hence, we test the performance of different LMMs (Large Multimodal Models), including GPT4-o, GPT-4, and InternVL2 (Chen et al., 2024a). Table 5 shows the ablation results on the figure data. To our surprise, the figure data doesn’t improve the LMMs’ results in this task, even harming the performances. This might be due to the low informativeness of the figures, as figures usually consume more input tokens but act only as supplementary information to the text, indicating future work on developing LMMs that can effectively leverage the scientific figures.
5.3 PaperWeakness
Settings.
Intuitively, the full paper content is necessary for providing feedback to that paper. Therefore, instead of setting a maximum input length, in Weakness, we try to feed all the paper context into the LLMs. As the input length of Weakness is extremely long (see Table 13), we adopt a “split-combine” method — we first split the whole paper into several smaller pieces and let LLMs predict the weaknesses of each piece separately; after that, we combine all pieces’ weaknesses as a final complete prediction. In practice, for the length of each small piece, we set 2,000 and 3,000 words for open- and closed-source LLMs, respectively. Additionally, in this task, we also examine the performance of a recent agent framework, namely AI-SCI (Lu et al., 2024), which enhances GPT-4o’s paper review ability by leveraging advanced prompting techniques, e.g., self-reflection (Shinn et al., 2024) and response ensembling (Wang et al., 2023). [^4] We don’t run AI-SCI on ExpDesign, because AI-SCI takes model-generated ideas as the inputs, which are incompatible with our task setting.
Main results.
Table 6: Various LLMs’ performances on the 993 instances of Weakness.
Weakness Diversity
Methods SN-F1 (%) SN-Precision (%) SN-Recall (%) ITF-IDF ($\uparrow$)
Human Review 7.69
Open-source LLMs
OLMo-7B (Groeneveld et al., 2024) 43.25 40.38 47.04 2.45
Falcon-40B (Almazrouei et al., 2023) 27.34 25.13 30.88 1.06
Gemma 2-27B (Gemma Team, 2024) 35.85 34.68 37.91 1.43
Mistral-7B (Jiang et al., 2023) 42.03 43.80 40.77 1.17
Mixtral-8x22B-MoE (Jiang et al., 2024) 43.23 44.59 42.23 0.98
Llama 3.1-70B (MetaAI, 2024) 42.78 43.19 42.70 2.60
Qwen 2.5-72B (Qwen Team, 2024) 42.74 43.80 42.05 1.21
Closed-source LLMs
Gemini 1.5 Pro (Anil et al., 2023) 48.75 43.97 55.08 5.88
Claude 3.5 sonnet (Anthropic, 2024a) 47.85 41.97 56.00 3.91
GPT-4 (OpenAI et al., 2023) 47.66 42.15 55.19 5.31
GPT-4o (OpenAI, 2024a) 47.73 42.09 55.48 5.95
o1-preview (OpenAI, 2024b) 48.62 42.54 57.08 5.63
LLM Agent Framework
AI-SCI (GPT-4o) (Lu et al., 2024) 45.05 40.02 51.91 2.23
Table 6 shows the main results, where the closed-source LLMs’ overall performances are generally superior to the results of open-source LLMs. Similarly, closed-source LLMs are particularly excellent in SN-Recall because of more generated weaknesses. However, there is still a considerable gap in the weakness diversity between the LLMs and human experts. [^5] Note that the human’s ITF-IDF score in Table 6 can be slightly underestimated. This is because we keep the repeated weaknesses in the human review, which affects the human review’s informativeness (lower ITF) but is useful when calculating the SN-Recall for LLMs. Compared with human review, most LLM-generated weaknesses are vague and lack the necessary knowledge about some frontier research works. Surprisingly, AI-SCI performs worse than backbone GPT-4o, especially on ITF-IDF, which suggests the challenge of Weakness, i.e., simply adopting popular prompting techniques cannot well address this task.
Table 7: The performance comparison of different input processing methods for Weakness. We use GPT-4o and GPT-4-Turbo because both accept a maximum of 128k tokens input. We also put the results of AI-SCI in the table for reference. Here, “split-combine” splits the input paper into several pieces, where each piece’s length is denoted as “window size”; “no-split” means the conventional input cutting, for example, if the window size is 3,000, then only the first 3,000 words in the paper are used. According to the data statistics, 20,000 words can cover maximum lengths of more than 95% of the papers in our dataset.
Models Input Context Processing Window Size (in words) SN-F1 SN-Precision SN-Recall ITF-IDF
GPT-4-Turbo split-combine 3,000 47.66 42.15 55.19 5.31
no-split 3,000 45.80 43.66 48.39 5.58
no-split 20,000 44.99 42.64 47.82 5.58
GPT-4o split-combine 3,000 47.73 42.09 55.48 5.95
no-split 3,000 45.74 43.45 48.54 5.92
no-split 20,000 45.47 42.97 48.51 6.02
AI-SCI split-combine 3,000 45.05 40.02 51.91 2.23
no-split 3,000 42.56 40.90 44.65 2.53
no-split 20,000 42.53 40.75 44.78 2.58
$\mathcal{Q}_{1}$: Is the split-combine effective?
Ideally, if the LLM has a sufficient context window size, it is unnecessary to split the input papers for separate processing. Consequently, in this paragraph, we utilize the LLMs accepting long context input to compare “split-combine” with “no-split”, i.e., letting LLMs write weaknesses by giving the full paper. In practice, we set the maximum number of input words to 20k, which ensures $\geq$95papers in the Weakness can be fully processed. As shown in Table 7, compared with giving the full paper contexts, split-combine generally brings about superior performances. During manual checking, we find that, when full paper is available, LLMs frequently neglect some important sections and omit weaknesses accordingly, while split-combine ensures that the LLMs can carefully brainstorm weaknesses within each smaller piece. Surprisingly, the LLMs’ performances with full paper context can be even worse than just remaining the first 3,000 words. This implies that even the current powerful long-context LLMs still fall short when processing long scientific documents.
$\mathcal{Q}_{2}$: Does multi-modal input boost performance?
Table 8: The ablation study about the paper tables and figures of Weakness. Based on the conclusion in Table 7, we use the “split-combine” to process the text input here (2,000 and 3,000 words context window size for open- and closed-source models). For GPT-4o, we use all the table/figure images; while for InternVL2, we randomly select two images per paper, i.e., two random figures, two random tables, or one random figure + table.
Models SN-F1 SN-Precision SN-Recall ITF-IDF
GPT-4o 47.73 42.09 55.48 5.95
w/ tables 46.76 41.32 54.17 5.53
w/ figures 46.62 41.20 54.04 5.48
w/ tables & figures 46.58 41.17 53.98 5.36
InternVL2-26B 41.91 41.02 43.28 1.48
w/ tables 40.55 40.37 42.91 1.46
w/ figures 42.88 42.10 43.76 1.46
w/ tables & figures 42.44 42.00 43.31 1.44
Our dataset covers both tables and figure illustrations extracted from the paper PDF as inputs. Intuitively, when reviewing a paper, both figures and tables are critical, not only for a better understanding, but also because some weaknesses are related to tables/figures. [^6] We find that there is approximately one human-written weakness related to figures or tables in each paper. Therefore, in Table 8, we adopt two LMMs to investigate the effectiveness of image inputs. Overall, image information, including both figures and tables, doesn’t bring significant performance improvement, i.e., only InternVL2 gains a performance boost after incorporating figures; while tables slightly drop both models’ results. This is probably because the LMMs cannot reason well over the information-intensive images, especially the table images.
5.4 ReviewCritique
Settings.
As individual review comments are split into multiple smaller segments (sentences), in order to avoid the performance variance that comes from the prompting, we follow Du et al. (2024) to utilize two prompting strategies. i) Labeling-All: given everything necessary including a list of indexed review segments, require the LLM to output a list of triples, like {id, reliable or not, explanation}. ii) Select-Deficient: Given everything necessary including a list of indexed review segments, require the LLM to output a list of tuples, {id, explanation}, when it believes the “id” corresponds to a deficient segment.
To further enhance evaluation robustness, we ensemble the results obtained from the two prompting strategies using two methods. i) Both “No”: if both prompts classify a segment as deficient, we consider it to be deficient. ii) Either “No”: if either of the prompts labels a segment as Deficient, we consider it to be deficient.
Main results.
Table 9: From (Du et al., 2024), various LLMs’ performances on the 11,376 instances of ReviewCritique. The best F1 score among different prompt methods for a single model is underlined. The best F1 score across all models is also bold.
Models Precision / Recall / F1
Labeling-All Select-Deficient Both “No” Either “No”
Open-source LLMs
Llama3-8B (AI@Meta, 2024) 7.73 / 45.95 / 12.22 11.47 / 30.29 / 14.88 11.37 / 21.27 / 12.46 8.19 / 53.61 / 13.35
Llama3-70B (AI@Meta, 2024) 13.63 / 42.49 / 18.19 13.95 / 31.16 / 17.46 16.16 / 23.51 / 16.67 12.46 / 50.02 / 18.43
Qwen2-72B (Bai et al., 2023) 9.97 / 26.60 / 12.96 11.35 / 34.61 / 14.64 9.07 / 15.13 / 9.62 10.49 / 43.00 / 15.16
Closed-source LLMs
Gemini 1.5 (Anil et al., 2023) 16.58 / 34.13 / 19.76 14.71 / 43.60 / 19.72 17.01 / 27.05 / 18.28 14.46 / 50.37 / 20.34
GPT-4 (OpenAI et al., 2023) 14.91 / 34.49 / 18.38 17.18 / 34.59 / 20.30 18.71 / 21.40 / 16.85 14.72 / 47.68 / 20.66
Claude Opus (Anthropic, 2024b) 16.86 / 34.26 / 20.35 17.69 / 26.61 / 18.71 17.14 / 18.70 / 15.78 16.94 / 42.12 / 21.99
We put the results of Du et al. (2024) in Table 9. Closed-source models (GPT-4, Claude Opus, and Gemini 1.5) generally outperform open-source models (Llama3-8B and 70B, Qwen2-72B) in F1 score. Claude Opus achieves the highest F1 scores, with GPT-4 and Gemini 1.5 performing slightly worse. Notably, recall scores are consistently higher than precision scores across all LLMs and prompting strategies, suggesting that LLMs tend to incorrectly identify segments as deficient. Despite the superior performance of the closed-source models, their F1 scores remain relatively low even with different prompt strategies, highlighting the challenges LLMs face in such expertise-intensive tasks and emphasizing the importance of human expertise in the meta-reviewing process.
$\mathcal{Q}$: How about the LLMs explanation quality regarding the deficient review?
As we also prompt the LLMs to generate the corresponding explanation on why they think each review segment is deficient, we report how well the model-generated deficiency explanation aligns with the human explanation.
Table 10: Evaluation of LLMs’ explanations for correctly identified deficient segments.
Model ROUGE-1/2/L/BERTScore
GPT-4 17.13 / 2.71 / 14.64 / 55.63
Claude Opus 20.18 / 3.69 / 17.52 / 57.28
Gemini 1.5 18.47 / 2.98 / 16.38 / 56.46
Llama3-8B 16.49 / 2.22 / 13.65 / 55.23
Llama3-70B 15.94 / 1.95 / 13.78 / 57.09
Qwen2-72B 17.07 / 3.00 / 14.69 / 56.88
We put the results of Du et al. (2024) in Table 10. The results in Table 10 show that overall scores for all LLMs are relatively low, indicating they can identify some Deficient segments but struggle to articulate their reasoning. Among the LLMs, Claude Opus achieves the highest scores across all metrics, suggesting its explanations align best with human annotators. Claude Opus also excels in identifying Deficient segments, as shown previously. GPT-4 and Gemini 1.5 show similar performance to Claude Opus. The open-source models, Llama3 (8B and 70B) and Qwen2-72B, generally score lower than the closed-source models.
6 Conclusion
In this work, we propose AAAR-1.0, a novel benchmark targeting a comprehensive evaluation of the current LLMs’ AI research capacity. AAAR-1.0 consists of distinct expertise-intensive tasks along with the curated evaluation metrics, and collect high-quality data by employing senior AI researchers. Extensive experiments highlight the challenges and values of AAAR-1.0.
Limitations
We shed light on two limitations of this work: i) As we gather data from open-source platforms such as arXiv and OpenReview, there is a possibility that current or future LLMs may be trained on the same source data utilized in our benchmark. This situation could influence the fairness of LLM comparisons and the conclusions drawn from this paper. While we acknowledge this potential data leakage, we maintain that our work can provide valuable insights and serve as an upper bound for some LLMs, particularly if they are indeed pretrained on those scientific papers. At present, this research can inform the design of current LLMs and agents, enhancing the community’s understanding of the strengths and limitations of using LLMs in scientific research. ii) Meanwhile, due to the high consumption of employing senior researchers in conducting data annotation, the data size for some tasks, such as ExpDesign, is relatively small. This might introduce variability in LLM performance, even with repeated runs. As this work marks the beginning of a series, we plan to release larger-scale datasets that will cover more diverse research tasks in AAAR-2.0.
Ethics Statement
Our study highlights that LLMs can be used to assist humans in AI research. This doesn’t mean we encourage the use of AI to replace human researchers. In contrast, we emphasize that the main responsibility for conducting scientific research should always lie with humans to prevent any societal risks, while the LLMs are only tools for making human research more efficient. To this end, we commit to careful distribution of the data collected in our research, ensuring it serves strictly for research purposes. Our goal is to mitigate the risks while maximizing the benefits offered by LLMs.
Acknowledgments
We would like to thank arXiv and OpenReview for releasing the paper source packages and the review comments. The authors would also like to thank Ibraheem Moosa and Sarkar Snigdha Sarathi Das for assisting in the data collection.
Appendices
Within this supplementary material, we elaborate on the following aspects:
  • Appendix A: Data Statistics and Diversity
  • Appendix B: Implementation Details
  • Appendix C: More Experiment Results and Details
  • Appendix D: Data Cases and Annotation Platform Illustration
  • Appendix E: Prompt Templates
Appendix A Data Statistics and Diversity
We provide the detailed data statistics of three datasets in our benchmark, as shown in Table 11, 12, and 13. We use the NLTK package [^7] https://www.nltk.org/ to tokenize words and count the length. When calculating the length of equations, we use the pylatexenc tool [^8] https://github.com/phfaist/pylatexenc to simplify the equations first.
Meanwhile, for the Weakness, we also plot the review scores distribution of the papers used in the dataset, as well as the track distribution. As can be found in Figure 5, our dataset has a decent distribution, where the papers are uniformly distributed across 13 tracks, and most papers’ scores ranged from 5 to 8 (i.e., most papers are weakly rejected or accepted).
[…]
# of classification instances 1,049
# of source papers 869
ave. “left” input context length (in words) 4,377
ave. “right” input context length (in words) 6,362
max “left” input context length (in words) 24,849
max “right” input context length (in words) 32,948
min “left” input context length (in words) 711
min “right” input context length (in words) 8
ave. “pos.” output equation length (in character) 55
ave. “neg.” output equation length (in character) 48
max “pos.” output equation length (in character) 1,039
max “neg.” output equation length (in character) 306
min “pos.” output equation length (in character) 6
min “neg.” output equation length (in character) 4
Table 12: The statistics of ExpDesign.
# of instances 100
# of source papers 100
ave. input context length (in words) 4,288
max input context length (in words) 9,799
min input context length (in words) 698
ave. # of input figures 2.6
max # of input figures 16.0
min # of input figures 0.0
ave. length of Experiment&Explanation list 5.7
ave. length per experiment (in words) 34.3
ave. length per explanation (in words) 27.1
max length of Experiment&Explanation list 13
max length per experiment (in words) 135
max length per explanation (in words) 89
min length of Experiment&Explanation list 2
min length per experiment (in words) 9
min length per explanation (in words) 9
Table 13: The statistics of Weakness.
# of instances 993
# of source papers 993
ave. input context length (in words) 9,811
max input context length (in words) 49,195
min input context length (in words) 24
ave. # of input figures 7.0
max # of input figures 37.0
min # of input figures 0.0
ave. # of input tables 4.3
max # of input tables 53.0
min # of input tables 0.0
ave. # of reviewers per paper 3.8
max # of reviewers per paper 9.0
min # of reviewers per paper 3.0
ave. # of weaknesses per reviewer 4.8
max # of weaknesses per reviewer 39.0
min # of weaknesses per reviewer 1.0
ave. length of weakness (in words) 39.1
max length of weakness (in words) 371.0
min length of weakness (in words) 2.0
(a) The review score distribution of the papers used in Weakness.
(b) The track distribution of the papers used in Weakness.
Figure 5: The data diversity illustration of Weakness, including the score distribution and track distribution of the papers used in our dataset.
Appendix B Implementation Details
B.1 Metric Details
When calculating the metrics, specifically for the similarity-based scores, we utilize SentenceBERT (Reimers, 2019) to encode each segment (e.g., each experiment idea in the list) into a dense vector, and then calculate the cosine similarity, [^9] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 which takes about 1GB of memory when running on a single A100 GPU.
B.2 LLMs Running Details
In our experiments, we utilize various LLMs, including both closed and open-sourced. We list the model weight sources for the open-source LLMs:
We use VLLM to unify the inference endpoints of all the above models. [^10] https://github.com/vllm-project/vllm We use Pytorch 2.4.0 with CUDA 12.1, and use 8 NVIDIA A100 GPUs for the LLMs inference.
Meanwhile, we use the gpt-4o-2024-08-06, gpt-4-1106-preview, o1-preview-2024-09-12, gemini-1.5-pro-002, and claude-3-5-sonnet-20240620 for the closed-source LLMs. We use LiteLLM to unify the API calling for all these LLMs. [^11] https://github.com/BerriAI/litellm
Given the unstable performance of LLMs, particularly closed-source ones, we run each model thrice during our experiments, selecting the median result from these repeated runs.
Appendix C More Experiment Results and Details
C.1 Human Evaluation on LLM-Generated Explanation
We ask 5 annotators to evaluate the LLM-generated explanations. Specifically, each of them is assigned 4 or 5 papers, along with the corresponding experiment lists. For each paper, the annotator is given 5 different models’ outputs (model names are anonymized), and the annotator has to decide if each LLM-generated explanation is acceptable according to the experiment. We show the human evaluation results in Table 4,
Appendix D Data cases and Annotation Platform Illustration
As shown in Figure 7, 8, and 9, we show the sample cases of the three tasks in AAAR-1.0. Meanwhile, we illustrate the screenshot of our annotation platform in Figure 6.
Figure 6: The annotation platform for collecting the annotation of ExpDesign. We ask annotators to first make comments on the Google Drive PDF, then move all the annotations to the online Google Doc (for further verification and discussion).
Appendix E Prompt Templates
In this appendix, we attach all the prompts used in this work, including prompts in data collection and model prediction, as shown in Figure 10, 11, and 12.
Figure 7: A sample case of EqInfer.
Figure 8: A sample case of ExpDesign.
Figure 9: A sample case of Weakness.
Figure 10: The prompts used in EqInfer, including both data collection and model prediction.
Figure 11: The prompts used in ExpDesign, including both data collection and model prediction.
Figure 12: The prompts used in Weakness.

### v2

@@
-Assist Research
+Renze Lou    Hanzi Xu    Sijia Wang    Jiangshu Du    Ryo Kamoi    Xiaoxin Lu    Jian Xie    Yuxuan Sun    Yusen Zhang    Jihyun Janice Ahn    Hongchao Fang    Zhuoyang Zou    Wenchao Ma    Xi Li    Kai Zhang    Congying Xia    Lifu Huang    Wenpeng Yin
 Abstract
@@
 However, researchers face unique challenges and opportunities in leveraging LLMs for their own work, such as brainstorming research ideas, designing experiments, and writing or reviewing papers.
-In this study, we introduce AAAR-1.0, a benchmark dataset designed to evaluate LLM performance in four fundamental, expertise-intensive research tasks: (i) EquationInference, assessing the correctness of equations based on the contextual information in paper submissions; (ii) ExperimentDesign, designing experiments to validate research ideas and solutions; (iii) PaperWeakness, identifying weaknesses in paper submissions; and (iv) ReviewCritique, identifying each segment in human reviews is deficient or not.
+In this study, we introduce AAAR-1.0, a benchmark dataset designed to evaluate LLM performance in three fundamental, expertise-intensive research tasks: (i) EquationInference, assessing the correctness of equations based on the contextual information in paper submissions; (ii) ExperimentDesign, designing experiments to validate research ideas and solutions; (iii) PaperWeakness, identifying weaknesses in paper submissions; and (iv) ReviewCritique, identifying each segment in human reviews is deficient or not.
 AAAR-1.0 differs from prior benchmarks in two key ways: first, it is explicitly research-oriented, with tasks requiring deep domain expertise; second, it is researcher-oriented, mirroring the primary activities that researchers engage in on a daily basis.
 An evaluation of both open-source and closed-source LLMs reveals their potential as well as limitations in conducting sophisticated research tasks.
-We will release the AAAR-1.0 and keep iterating it to new versions.
+We will keep iterating AAAR-1.0 to new versions.
 Project Webpage:   https://renzelou.github.io/AAAR-1.0/
+Machine Learning, ICML
+[…]
+[…]
 Figure 1: The input-output illustration of four tasks in the proposed AAAR-1.0 benchmark.
@@
 To this end, in this work, we introduce AAAR-1.0, a novel benchmark that aims to comprehensively assess the LLMs’ capacity on expert-level research tasks.
-As illustrated in Figure 1, AAAR-1.0 decomposes four distinct expert-level AI research tasks from the researcher’s daily activities, including i) EquationInference, investigating whether the LLMs can infer the equation correctness based on the paper context; ii) ExperimentDesign, validating LLMs’ ability on designing reliable experiments for a research idea; iii) PaperWeakness, testing the quality of weaknesses discovered by LLMs from paper drafts; and iv) ReviewCritique, investigating whether LLMs can identify and explain the deficient/unreliable human-written paper reviews.
+As illustrated in Figure 1, AAAR-1.0 decomposes three distinct expert-level AI research tasks from the researcher’s daily activities, including i) EquationInference, investigating whether the LLMs can infer the equation correctness based on the paper context; ii) ExperimentDesign, validating LLMs’ ability on designing reliable experiments for a research idea; iii) PaperWeakness, testing the quality of weaknesses discovered by LLMs from paper drafts; and iv) ReviewCritique, investigating whether LLMs can identify and explain the deficient/unreliable human-written paper reviews.
 To ensure data quality, senior AI researchers with extensive domain expertise perform data annotation for AAAR-1.0, followed by rigorous multi-round data examination and filtering.
-All four tasks require models to possess strong domain knowledge covering various cutting-edge research findings, as well as expert-level research experience, to the extent that even humans need substantial research accumulation to tackle the tasks we designed.
+All three tasks require models to possess strong domain knowledge covering various cutting-edge research findings, as well as expert-level research experience, to the extent that even humans need substantial research accumulation to tackle the tasks we designed.
 Crucially, tasks here are singular, stand-alone challenges (with clear input and output expectations) rather than a complicated task chain (Li et al., 2024; Lu et al., 2024), providing a more transparent assessment of the model’s intermediate output.
 Benefiting from the proposed automatic metrics, we conduct extensive experiments across numerous mainstream LLMs, where we find that:
-  • With a random guess baseline of 25%, the performance of most LLMs on EqInfer hovers just slightly above chance, with the top models reaching around 60%.
-This highlights the difficulty of the task, despite its reliance only on local context reasoning.
+  • With a random guess baseline of 40% F1, the performance of most LLMs on EqInfer hovers just slightly above chance, with the top models reaching around 46%.
+This highlights the difficulty of the task, despite its reliance primarily on local context reasoning.
   • In ExpDesign, LLM-designed experiments are innovative and more diverse than those by humans; however, many are trivial, lack feasibility, and stray from the original research objectives.
@@
   • In ReviewCritique, LLMs struggle to effectively identify deficient human reviews, indicating limited usefulness in assisting meta-reviewers in evaluating the quality of individual paper reviews.
+LLMs generated weakness lacks sufficient domain-specific knowledge, model always tends to generate some vague weaknesses that can be applied to any other papers.
 2 Related Work
 LLMs for AI Research.
-With the rapid evolution of pertaining techniques, LLMs are found to be useful in assisting various research disciplines (Yu et al., 2024; Labrak et al., 2024), particularly in AI research, such as generating novel research ideas (Kumar et al., 2024), reviewing research draft (Gao et al., 2024; Du et al., 2024; Liang et al., 2024), and writing scientific papers (Chamoun et al., 2024; Lu et al., 2024).
+With the rapid evolution of pertaining techniques, LLMs are found to be useful in assisting various research disciplines (Yu et al., 2024a; Labrak et al., 2024), particularly in AI research, such as generating novel research ideas (Kumar et al., 2024; Yu et al., 2024b), reviewing research draft (Gao et al., 2024; Du et al., 2024; Liang et al., 2024), and writing scientific papers (Chamoun et al., 2024; Lu et al., 2024).
 For example, Si et al.
@@
 (2024) found that while LLMs are effective at summarizing papers, they tend to overly trust the authors’ claimed strengths and struggle to identify weaknesses specific to the paper.
-Furthermore, some works try to employ LLMs to solve more complicated research tasks that are composed of multiple steps (Li et al., 2024; 2023; Tang et al., 2023).
+Furthermore, some works try to employ LLMs to solve more complicated research tasks that are composed of multiple steps (Li et al., 2024, 2023; Tang et al., 2023).
 Notably, Lu et al.
@@
 Benchmarks for AI Research Tasks.
-Existing “LLM assists research” benchmarks mainly focus on the implementation and execution part of the research pipeline (Lu et al., 2024; Chen et al., 2024b; Li et al., 2024).
+Existing “LLM assists research” benchmarks mainly focus on the implementation and execution part of the research pipeline (Lu et al., 2024; Chen et al., 2024a; Li et al., 2024; Chan et al., 2024).
 For instance, Huang et al.
 (2024) proposed MLAgentBench to test the LLMs’ capacity for writing project code and training the ML models, where the evaluation metric is the test performance of the models trained by LLMs.
-However, real-world AI research activities are diverse and some of them are hard to assess for quality, such as generating research ideas, which requires intensive manual assessment (Si et al., 2024; Liang et al., 2024), or LLM-based estimation (Lu et al., 2024).
+However, real-world AI research activities are diverse and some of them are hard to assess for quality, such as generating research ideas, which requires intensive manual assessment (Si et al., 2024; Liang et al., 2024).
 Our work centers on tasks that emphasize a comprehensive mastery of the scientific research field and core elements of a researcher’s daily workload, and we try to build curated task-specific metrics for every single task for a more efficient and accurate LLMs appraisal.
@@
 3 AAAR-1.0
-Figure 2 provides an overview of the construction process of AAAR-1.0.
-In the following sections, we elaborate on the data collection details of the aforementioned four tasks, including § 3.1 EquationInference (EqInfer), § 3.2 ExperimentDesign (ExpDesign), § 3.3 PaperWeakness (Weakness), and § 3.4 ReviewCritique.
+Figure 2 provides a data construction overview.
+In the following sections, we elaborate on the data collection details, including § 3.1 EquationInference (EqInfer), § 3.2 ExperimentDesign (ExpDesign), and § 3.3 PaperWeakness (Weakness), and § 3.4 ReviewCritique.
 3.1 EquationInference
@@
 Directly prompting LLMs to generate equations proves overly demanding.
-Therefore, this work formulates EqInfer (Figure 1) as as a multiple-choice equation inference task based on contextual cues, curated as follows.
+Therefore, this work formulates EqInfer (Figure 1) as a binary inference task.
+[^1] EqInfer also facilitates a multiple-choice QA setting; while we find a binary inference is more challenging for LLMs.
 ① Data crawling and cleaning.
 For the data source, we adopt the pre-compilation LaTeX code for two reasons: i) existing PDF parsing tools, such as PyMuPDF and PaperMage (Lo et al., 2023), can introduce considerable noise to the parsed equation text; ii) considering most of exiting LLMs are capable with processing LaTeX code, using LaTeX source instead of parsed text can be more accurate and provide LLMs with richer information.
-Meanwhile, to avoid using any low-quality human-written equations, we only crawl those peer-reviewed papers accepted by top-tier conferences.
+Meanwhile, we only crawl those peer-reviewed papers accepted by top-tier conferences to avoid using low-quality human-written equations.
 Accordingly, we first obtain the accepted paper list from ACL Anthology, from year 2019 to 2023.
@@
 We then clean the LaTeX sources by deleting all the comments and combining multiple cross-referred .tex files into a main file.
-Afterwards, we use regex to randomly extract (at most) 3 equations’ code snippets per paper, finally resulting in 3,877 human-written equations are extracted.
+Afterward, we use regex to randomly extract (at most) 3 equations’ code snippets per paper, resulting in 3,877 human-written equations.
 ② LLM-based equation synthesis.
-As we formulate this task as classification, for each human-written positive equation, we have to craft at least three counterpart negative equations.
-To this end, we prompt GPT-4 to synthesize more equations based on the paper context.
-For each positive equation, we repeat this prompting (with a high decoding temperature) until three different negative equations are synthesized.
+As EqInfer assessing whether the LLMs can infer the correctness of equation (i.e., binary classification), for each human-written positive equation, we have to craft counterpart negative equations.
+To this end, for each positive equation, we prompt GPT-4 to synthesize a negative equation based on the paper context.
+We repeat this prompt (with a high decoding temperature) until three different negative equations are synthesized.
+[^2] The number of negative equations is empirically decided.
 ③ LLM-based filtering.
-However, the LLM-synthetic equations can sometimes be context-unaligned, i.e., some synthesized equations contain notations that are never defined in the paper context, which is a superficial shortcut for the classification tasks.
-To improve the data quality, we prompt GPT-4 to identify those context-unaligned negative equations.
-We then discard those instances where all three negative equations are identified as contextually unaligned.
-This filtering leads to a final of 1,449 classification instances (62.3% instances are filtered).
+However, the LLM-synthetic equations can be context-unaligned, i.e., some synthesized equations contain notation that is never defined in the paper context, which becomes a superficial shortcut and too effortless for LLMs to identify.
+To improve data quality, we prompt GPT-4 to identify context-unaligned negative equations.
+We then eliminate the positive equation and its negative counterparts, where all three negative counterparts are unaligned.
+This filtering leads to a final of 1,449 positive equations and 4,347 negative equations (each positive equation has three negative counterparts, and at least one negative counterpart is “challenging”).
 ④ Expert-based examination.
-Furthermore, it’s also possible that synthesized negative equations are actually correct (i.e., false negative options) — even if the negative and positive equations are written differently, the final compiled results might be the same.
-To filter out the false negative equations and to have a final check on the classification instances, we then employ human experts to conduct a further data review.
-We asked 5 senior Ph.D.
-students who are experienced in AI research to manually check all the instances.
-For each classification instance, we ask those human experts to consider the following criteria: i) Are all four equations (both positive and negative) grammatically correct?
-ii) After compilation, is there only one correct answer?
-We ask every human expert to use external LaTeX compilation tools (e.g., TeXlive), and identify the instances that cannot meet the criteria.
-Each instance is examined by at least two experts, and we only keep instances that all experts decide to keep.
-After this strict examination, a total of 1,049 instances are eventually kept (27.6% instances are filtered)
+Furthermore, it’s also possible that synthesized negative equations are actually correct (i.e., false negative) — even if the negative and positive equations are written differently, the final compiled results might be the same.
+We then employ human experts to review the data further and filter out false negative equations, checking the classification instances for accuracy.
+We asked 5 senior PhD students who are experienced in AI research to check all instances.
+We ask human experts to consider the following criteria for each positive equation and its negative counterparts (each pair): i) Are all equations grammatically correct?
+ii) After compilation, are all negative equations different from the positive ones?
+We ask every human expert to use external LaTeX compilation tools (e.g., TeXlive), and identify the pairs that cannot meet the criteria.
+Each pair is examined by at least two experts, and we only keep pairs that all experts decide to keep.
+After this strict examination, a total of 1,049 pairs are eventually kept (27.6% pairs are filtered)
 Final data.
-We finally shuffle the four equations for each classification instance and randomly assign letters (A, B, C, and D) to the equations.
-We show the data statistics of the final EqInfer in Table 11 and the sample data cases in Figure 7.
+We finally obtain 1,049 positive equations (each has three negative counterparts). We show data statistics of  EqInfer in Table 9 and data examples in Figure 8.
 3.2 ExperimentDesign
@@
 Hence, we conduct a further multi-round peer discussion.
-For each online paper PDF, where all the key experiments are identified, summarized, and explained, we ask a different expert (reviewer) to review the annotation by considering the following three criteria: i) Are the identified experiments all the key experiments?
+For each paper, where all the key experiments are identified, summarized, and explained, we ask a different expert (reviewer) to review the annotation by considering the following three criteria: i) Are the identified experiments all the key experiments?
 ii) Does each experiment summarization covers all key information?
 iii) Does each explanation sound reasonable and reliable?
-Each reviewer has to leave comments to the online PDF regarding the above criteria, and then the annotator has to respond to each comment — either accept the suggestion and revise the previous annotation, or provide a “rebuttal” to the reviewer to uphold the annotation.
-This discussion iterates until both opinions align with each other.
+Each reviewer must leave comments on the online PDF regarding the above criteria, and then the annotator must respond to each comment — either accept the suggestion and revise the previous annotation or provide a “rebuttal” to the reviewer to uphold the annotation.
+This discussion is iterative until both opinions align.
 Eventually, for each paper, we collect two lists: i) the experiment list, summarizing each experiment step of the paper; ii) the explanation list, the underlying motivations that are one-one corresponding to the experiment.
@@
 4 Evaluation Criteria
-For EqInfer, we adopt accuracy as the classification criterion.
-For ExpDesign and Weakness, since both tasks have natural language outputs, semantic-based metrics are necessary.
-Hence, in addition to the conventional ROUGE (Lin, 2004), we also develop several novel similarity-based metrics for each specific task.
-When evaluating the experiment plan list of ExpDesign, we hope the LLMs can mention as many similar experiment steps as the expert’s plan.
-Nevertheless, we also don’t expect LLMs to generate too many irrelevant or redundant steps in the plan.
-This intuition covers both the “recall” and “precision” aspects.
-Therefore, we develop semantic similarity-based F1 score, denoted as S-F1, which is the harmonic mean of S-Precision and S-Recall:
-S-Precision $=\frac{1}{m}\sum_{i=1}^{m}\max_{j}\textrm{sim}(p_{i},g_{j})$ (1)
-S-Recall $=\frac{1}{n}\sum_{j=1}^{n}\max_{i}\textrm{sim}(g_{j},p_{i})$ (2)
-where the $p$ and $g$ represent the LLM’s prediction plan and the ground-truth plan, respectively.
-The $m$ and $n$ are the list length of $p$ and $g$ (e.g., $m$ experiment steps in $p$).
-We use SentenceBERT (Reimers, 2019) to measure the semantic similarity between the $p_{i}$ step and the $g_{j}$ step.
-Meanwhile, S-F1 omits the item order difference of two lists, but when giving same-length lists (items have one-one correspondence), we can utilize the following similarity-based matching score:
+For EqInfer, we adopt F1 as the classification criterion.
+For ExpDesign and Weakness, since both tasks have free-form outputs, we develop several novel task-specific metrics in addition to the conventional ROUGE (Lin, 2004).
+We use LLMs to evaluate the experiment list of ExpDesign. Specifically, given a model-predicted experiment list $p$, and the ground-truth list $g$, we calculate:
+En-Precision $=\frac{1}{m}\sum_{i=1}^{m}f(p_{i},g)$ (1)
+En-Recall $=\frac{1}{n}\sum_{j=1}^{n}f(g_{j},p)$ (2)
+where the $m$ and $n$ are the list length of $p$ and $g$; $f(.)$ represents the LLM prompting, where we prompt LLM to decide whether each predicted experiment item ($p_{i}$) is entailed by the whole ground-truth list ($g$), proceeding with binary output, and vice versa.
+Intuitively, En-Precision reflects how many prediction experiments match ground-truth experiments.
+In this work, we used GPT-4o as an evaluator.
+While for the explanation generation of ExpDesign, as the prediction experiments are one-on-one corresponding to the ground truth, we adopt a semantic-based metric:
 […]
-Unlike ExpDesign, the output of Weakness is multiple reviewers’ weakness lists, which means we have to measure LLM’s single prediction list with a “nested” list.
-Hence, we rewrite S-Precision ,  S-Recall to SN-Precision , SN-Recall:
-SN-Precision $=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{r}\sum_{k=1}^{r}\max_{j}\textrm{sim}(p_{i},g^{k}_{j})\right)$ (4)
-SN-Recall $=\frac{1}{r}\sum_{k=1}^{r}\left(\frac{1}{n_{k}}\sum_{j=1}^{n_{k}}\max_{i}\textrm{sim}(g^{k}_{j},p_{i})\right)$ (5)
+where we use SentenceBERT (Reimers, 2019) to measure the semantic similarity between $p_{i}$ and $g_{j}$.
+Unlike ExpDesign, the ground truth of Weakness is multiple reviewers’ weakness lists.
+Instead of merely merging the opinions of various reviewers into one flattened list and keeping LLM-as-judge as the metric (which is not only costly but also loses the structural information of diverse research perspectives), we employ the following semantic-based metric to efficiently evaluate predicted weaknesses:
+S-Precision $=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{r}\sum_{k=1}^{r}\max_{j}\textrm{sim}(p_{i},g^{k}_{j})\right)$ (4)
+S-Recall $=\frac{1}{r}\sum_{k=1}^{r}\left(\frac{1}{n_{k}}\sum_{j=1}^{n_{k}}\max_{i}\textrm{sim}(g^{k}_{j},p_{i})\right)$ (5)
 where $r$ is the number of reviewers of the given paper, $n_{k}$ means the length of $k$-th reviewer’s weakness list, and $g^{k}_{j}$ indicates the $j$-th item in $k$-th reviewer’s weakness list.
@@
 Meanwhile, we also hope the review is informative, i.e., no excessive similar weaknesses in one review.
-Therefore, we adopt the following ITF-IDF metric proposed by Du et al.
-(2024), which is inspired by the classic TF-IDF:
+Inspired by the classic TF-IDF, we propose a novel review diversity metric:
 ITF-IDF $=\frac{1}{w}\sum_{j=1}^{w}\left(\frac{1}{m_{j}}\sum_{i=1}^{m_{j}}\log\left(\frac{m_{j}}{O_{i}^{j}}\right)\times\log\left(\frac{w}{R_{i}^{j}}\right)\right)$ (6)
@@
 In this section, we conduct extensive experiments on AAAR-1.0, across various mainstream LLMs, to quantify the current LLMs’ capacity to tackle high-level research tasks.
-Specifically, § 5.1 for EqInfer, § 5.2 for ExpDesign, § 5.3 for Weakness, and § 5.4 for ReviewCritique.
+Specifically, § 5.1 for EqInfer, § 5.2 for ExpDesign, § 5.3 for Weakness; and § 5.4 for ReviewCritique.
 Please refer to Appendix B.2 for running details of the LLMs.
 5.1 EquationInference
-Table 1: Various LLMs’ performances on the 1,049 instances of EqInfer task.
-Methods Accuracy (%)
-Random Guess 25.00
+Table 1: Various LLMs’ performances on EqInfer task (1,049 positive and 3,147 negative samples). “All-positive” indicates a baseline that predicts all equations as positive.
+Methods F1 Prec. Rec.
+All-Positive 40.00 25.00 100.00
 Open-source LLMs
-Gemma 2-27B (Gemma Team, 2024) 3.24
-Falcon-40B (Almazrouei et al., 2023) 4.39
-OLMo-7B (Groeneveld et al., 2024) 19.00
-Mistral-7B (Jiang et al., 2023) 22.21
-Qwen 2.5-72B (Qwen Team, 2024) 35.93
-Mixtral-8x22B-MoE (Jiang et al., 2024) 37.08
-Llama 3.1-70B (MetaAI, 2024) 38.13
+OLMo-7B (Groeneveld et al., 2024) 13.64 11.93 15.91
+Mistral-7B (Jiang et al., 2023) 28.45 19.28 54.24
+Mixtral-8x22B-MoE (Jiang et al., 2024) 40.90 26.15 93.80
+Qwen 2.5-72B (Qwen Team, 2024) 31.22 26.28 57.40
+Llama 3.1-70B (MetaAI, 2024) 33.08 22.14 65.39
 Closed-source LLMs
-Gemini 1.5 Pro (Anil et al., 2023) 34.31
-GPT-4o (OpenAI, 2024a) 43.18
-GPT-4 (OpenAI et al., 2023) 49.85
-o1-preview (OpenAI, 2024b) 59.49
-Claude 3.5 sonnet (Anthropic, 2024a) 61.10
-Figure 3: The input context length scaling trend on the EqInfer task.
+Gemini 1.5 Pro (Anil et al., 2023) 46.74 32.05 86.27
+Claude 3.5 sonnet (Anthropic, 2024a) 45.13 29.48 96.18
+GPT-4o (OpenAI, 2024a) 40.35 30.79 58.53
+o1-preview (OpenAI, 2024b) 46.35 31.43 88.27
 Settings.
 As different LLMs have distinct context windows, to ensure a fair comparison, we fix the maximum input length for all models.
-According to the data statistics of Table 11, we empirically use 1,000 words for both contexts before and after equations, i.e., 2,000 surrounded words.
+According to Table 9, we empirically use 1,000 words for both contexts before and after equations, i.e., 2,000 surrounding words.
 Main results.
 Table 1 shows the main results.
-Firstly, the open-source LLMs, especially the Falcon and Gemma, perform unexpectedly disappointing (even worse than random guesses).
-These screwed scores are mainly due to the poor long-context instruction following ability, where we find some open-source LLMs are confused with the massive input and often copy the LaTeX code from the input.
-In contrast, closed-source LLMs generally achieve superior accuracy, probably owing to the richer scientific knowledge from the larger model parameters.
-However, considering the conventional multi-choice QA formulation of EqInfer, the recently-released GPT-4o solely gets 43.18, implying the unique challenge of EqInfer compared with other scientific QA benchmarks (Song et al., 2023).
-Notably, with the help of internal CoT, o1 gains stronger performances than GPT-4/GPT-4o, indicating the potential benefits of adopting reasoning for this task.
+Firstly, a simple baseline that predicts all equations as positive achieves 40% F1 (due to the 1:3 of positive and negative equations), while nearly all open-source LLMs even cannot beat this naive baseline.
+Notably, though the performance of Mixtral is slightly superior to the baseline, the extremely biased precision and recall scores imply that Mixtral is also simply predicting almost all samples as positive instead of truly inferring.
+Meanwhile, compared to the All-Positive baseline, the performance superiority of the strong close-source LLMs is not significant, the best LLM on this task only obtains 46.74%, which demonstrates the challenge of EqInfer compared with other similar benchmarks (Song et al., 2023).
+The generally high recall with low precision of all LLMs also indicates real-world risks, e.g., relying on LLMs to check the validity of equations in paper review.
 $\mathcal{Q}$: Do more contexts boost performance?
-Table 1 unifies the input context lengths to 1,000 tokens for various LLMs.
-To answer this question, we experiment with long-context LLMs to investigate the impact of the input context lengths.
-Particularly, we scale the input length (per side) from 100 to 1,500 words.
-As shown in Figure 3, for the open-source LLMs (Llama and Qwen), after 300 words length, increasing the input context doesn’t help the performance and even significantly drops Qwen’s scores.
-While for the closed-source GPT-4-Turbo and GPT-4o, scaling up input length gradually boosts the performances at the first 1,000 words, but stabilizes afterwards.
-This is in line with human intuition, i.e., surrounding context is required for the equation inference, as the adjacent context usually provides important information, such as the target algorithm description or the notation definition.
-However, after exceeding a specific threshold, more context information is not beneficial anymore and even confuses those LLMs with poor long-context handling capacity (Wang et al., 2024; Liu et al., 2024).
+EqInfer places high demands on reasoning within the scientific context.
+To quantify the impact of input context length, we scale the input length (per side) from 100 to 1,500 words.
+As shown in Figure 4, for the open-source LLMs (Llama and Qwen), an appropriate context length can boost the performance; while for GPT-4o, scaling up the context length doesn’t contribute much to the F1.
+However, during the scaling, we find that the precision of GPT-4o is gradually increased, and the recall is decreased accordingly; considering the label distribution of EqInfer, we believe precision can better reflect the model’s true capacities on this task.
+Thus, we anticipate that scaling up context shall be beneficial to those strong close-source LLMs such as GPT-4o.
 5.2 ExperimentDesign
@@
 The explanation generation is based on the oracle experiments to prevent error propagation.
-“Copy Input” is a random baseline: for experiment design, randomly select 5 sentences from the input paper; for experiment explanation, directly copy each experiment idea.
+“Copy Input” directly copies each experiment idea as the explanation.
 Experiment Design Experiment Explanation
-Methods S-F1 S-Precision S-Recall S-Match ROUGE-L ROUGE-1
-Copy Input 21.13 17.94 26.76 40.32 22.06 25.28
+Methods En-F1 En-Precision En-Recall S-Match ROUGE-L ROUGE-1
+Copy Input 40.32 22.06 25.28
 Open-source LLMs
-OLMo-7B (Groeneveld et al., 2024) 33.94 37.25 31.79 45.78 26.30 30.38
-Falcon-40B (Almazrouei et al., 2023) 17.87 21.78 15.35 17.03 12.10 12.72
-Gemma 2-27B (Gemma Team, 2024) 34.33 39.71 30.51 42.77 26.20 29.63
-Mistral-7B (Jiang et al., 2023) 37.62 43.09 34.19 50.18 30.20 34.69
-Mixtral-8x22B-MoE (Jiang et al., 2024) 42.21 50.13 36.82 49.07 29.96 34.53
-Llama 3.1-70B (MetaAI, 2024) 40.57 48.43 35.43 50.05 29.33 34.11
-Qwen 2.5-72B (Qwen Team, 2024) 43.24 51.73 37.55 51.12 29.46 34.68
+OLMo-7B (Groeneveld et al., 2024) 14.80 17.50 19.80 45.78 26.30 30.38
+Mistral-7B (Jiang et al., 2023) 18.96 24.83 21.38 50.18 30.20 34.69
+Mixtral-8x22B-MoE (Jiang et al., 2024) 23.16 24.45 30.57 49.07 29.96 34.53
+Llama 3.1-70B (MetaAI, 2024) 22.92 23.10 29.76 50.05 29.33 34.11
+Qwen 2.5-72B (Qwen Team, 2024) 24.28 22.48 34.44 51.12 29.46 34.68
 Closed-source LLMs
-Gemini 1.5 Pro (Anil et al., 2023) 51.87 50.77 53.37 52.87 28.52 33.80
-Claude 3.5 sonnet (Anthropic, 2024a) 48.74 46.49 51.53 53.03 18.75 26.15
-GPT-4 (OpenAI et al., 2023) 43.89 42.34 45.82 55.03 22.82 30.01
-GPT-4o (OpenAI, 2024a) 53.00 51.24 55.12 54.79 27.54 34.31
-o1-preview (OpenAI, 2024b) 46.67 45.04 48.70 58.55 29.11 36.70
+Gemini 1.5 Pro (Anil et al., 2023) 27.25 28.66 34.92 52.87 28.52 33.80
+Claude 3.5 sonnet (Anthropic, 2024a) 27.99 24.48 42.09 53.03 18.75 26.15
+GPT-4o (OpenAI, 2024a) 25.03 22.25 36.59 54.79 27.54 34.31
+o1-preview (OpenAI, 2024b) 30.13 28.13 38.59 58.55 29.11 36.70
 Settings.
@@
 Table 2 shows the main results.
-For the experiment design, the closed-source LLMs generally outperform open-source LLMs, and both closed-/open-source LLMs are superior to the “Copy Input” baseline (except the Falcon).
-Despite the higher S-Precision, the open-source LLMs are seriously deficient in S-Recall compared with closed-source LLMs ($\sim$10$\downarrow$).
-We find that closed-source LLMs are more creative in experiment design and tend to generate more experiment ideas than open-source LLMs (though most of the experiment ideas are trivial), leading to excellent S-Recall.
-As for the experiment explanation, the S-Match scores of closed-source LLMs still surpass the open-source LLMs, while the score difference is not significant.
-Furthermore, we find the negative correlation between S-Match and the ROUGE, where the ROUGE scores of closed-source LLMs are broadly inferior.
+For the experiment design, the closed-source LLMs generally outperform open-source LLMs.
+However, the score values of all LLMs are relatively low (20$\sim$30%), implying the LLMs consistently miss ground-truth experiments from the origin paper (low recall), and they tend to generate more novel experiments that didn’t show in the origin paper (low precision).
+As for the experiment explanation, the S-Match scores of closed-source LLMs still surpass the open-source LLMs.
+Furthermore, there is a negative correlation between S-Match and ROUGE score, where the ROUGE scores of closed-source LLMs are broadly inferior.
 We find that the open-source LLMs often try to copy the terms or phrases from the given experiment, or even simply paraphrase the experiment instead of explaining, which results in a high superficial overlap with the ground-truth explanation.
 This observation highlights the importance of adopting the proposed S-Match to avoid evaluation bias of traditional generation metrics.
-$\mathcal{Q}_{1}$: Can self-contained experiment design enhance the experiment explanation?
+$\mathcal{Q}_{1}$: What is the quality of the model-generated novel experiments?
+The low En-Precision of LLMs in Table 2 indicates the creativity of LLMs in generating novel experiments.
+We then randomly sample 15 papers from the ExpDesign and ask 3 experts to manually review the model-generated novel experiments.
+Specifically, we ask the experts to judge the necessity of the novel experiments, where we set three necessity levels: “A” indicates the experiment is necessary/mandatory to support the main claim, “B” represents optional/supplementary experiments, and “C” for those unrelated experiments (see Appendix C.2 for evaluation details).
+Table 3 shows the necessity scores of the three strongest LLMs.
+We find that LLMs consistently generate a lot of novel experiments, especially the Claude; though most of them are optional, even fancy/unrelated experiments, there are still a considerable amount of necessary experiments generated, e.g., the results of o1.
+We further find that some novel experiments can be regarded as useful supplementary analyses w.r.t.
+the human experiments.
+Table 13 shows examples of model-suggested experiments.
+Table 3: The human evaluation results on the novel experiments suggested by LLMs. “A”, “B”, and “C” represent the different quality level (i.e., necessity); “A” is the best level.
+Models # of novel EXP Necessity (%)
+A B
+Gemini 1.5 Pro 59 30.59 45.76
+Claude 3.5 sonnet 112 21.78 50.00
+o1-preview 71 35.84 36.61
+$\mathcal{Q}_{2}$: Can self-contained experiment design enhance the experiment explanation?
 Table 4: The impact on S-Match scores of maintaining the experiment’s self-containment for ExpDesign.
@@
 According to further manual checking, after maintaining the self-containment of the experiments, the LLMs can refer to other experiments and better grasp the underlying motivation of the current experiment.
-$\mathcal{Q}_{2}$: Do human evaluation results align with automatic metrics for explanation?
+$\mathcal{Q}_{3}$: Do human evaluation results align with automatic metrics for explanation?
 Table 5: The human evaluation results on LLMs’ output explanations of ExpDesign. “Acc. ratio” means how many model outputs are accepted by the annotator.
@@
 However, the performance ranking of both tables is perfectly correlated with each other (Spearman’s rank correlation coefficient $=$ 1), demonstrating the effectiveness of S-Match.
-$\mathcal{Q}_{3}$: Do more contexts boost performance?
-Figure 4: The input context length scaling trend of different LLMs on the ExpDesign task.
-We also investigate the impact of input context length for ExpDesign.
-As shown in Figure 4, we scale up the input pre-experiment context length from 0.1k to 10k tokens (10k tokens is the maximum paper context length in the dataset).
-For the experiment planning, more input context does improve the performance of different LLMs, while this benefit stops after exceeding 5k tokens, which is similar to EqInfer’s scaling results — after the necessary information has been covered, scaling more up doesn’t boost the performance.
-Meanwhile, the results of the experiment explanation demonstrate that explaining motivations almost doesn’t require any paper context, i.e., the LLMs solely rely on the given experiments.
-However, we do not expect this because we hope LLMs can explain the motivation based on a thorough understanding of the paper, just like how human experts do.
-Hence, there is still a considerable gap between the LLMs and humans in terms of grasping research motivations.
-$\mathcal{Q}_{4}$: Does multi-modal input boost performance?
-Table 5: The figure inputs ablation of ExpDesign.
-For the maximum text input length, same as the setting in Table 2, we use 2,000 and 3,000 words for open- and closed-source models, respectively.
-For the closed-source GPT-4o and GPT-4, as they have long context window sizes, we use all the figures of each paper.
-While for InternVL2, we randomly select two figures per input paper.
-Experiment Design Experiment Explanation
-Models S-F1 S-Precision S-Recall S-Match ROUGE-L ROUGE-1
-GPT-4o 53.00 51.24 55.12 58.54 29.25 35.50
-w/ figures 50.11 48.94 51.59 58.53 27.87 34.30
-GPT-4 43.89 42.34 45.82 56.95 25.98 33.37
-w/ figures 43.54 42.56 44.85 55.03 22.82 30.01
-InternVL2-26B 40.52 48.95 35.20 50.03 29.13 34.26
-w/ figures 38.83 46.91 33.70 50.29 29.29 34.06
-Intuitively, besides the text, when designing experiments for a given research topic, the figures can provide rich supplementary information, such as an algorithm illustration that can help better understand this research topic and underlying motivations.
-Hence, we test the performance of different LMMs (Large Multimodal Models), including GPT4-o, GPT-4, and InternVL2 (Chen et al., 2024a).
-Table 5 shows the ablation results on the figure data.
-To our surprise, the figure data doesn’t improve the LMMs’ results in this task, even harming the performances.
-This might be due to the low informativeness of the figures, as figures usually consume more input tokens but act only as supplementary information to the text, indicating future work on developing LMMs that can effectively leverage the scientific figures.
-5.3 PaperWeakness
-Settings.
-Intuitively, the full paper content is necessary for providing feedback to that paper.
-Therefore, instead of setting a maximum input length, in Weakness, we try to feed all the paper context into the LLMs.
-As the input length of Weakness is extremely long (see Table 13), we adopt a “split-combine” method — we first split the whole paper into several smaller pieces and let LLMs predict the weaknesses of each piece separately; after that, we combine all pieces’ weaknesses as a final complete prediction.
-In practice, for the length of each small piece, we set 2,000 and 3,000 words for open- and closed-source LLMs, respectively.
-Additionally, in this task, we also examine the performance of a recent agent framework, namely AI-SCI (Lu et al., 2024), which enhances GPT-4o’s paper review ability by leveraging advanced prompting techniques, e.g., self-reflection (Shinn et al., 2024) and response ensembling (Wang et al., 2023).
-[^4] We don’t run AI-SCI on ExpDesign, because AI-SCI takes model-generated ideas as the inputs, which are incompatible with our task setting.
-Main results.
 Table 6: Various LLMs’ performances on the 993 instances of Weakness.
 Weakness Diversity
-Methods SN-F1 (%) SN-Precision (%) SN-Recall (%) ITF-IDF ($\uparrow$)
+Methods S-F1 (%) S-Precision (%) S-Recall (%) ITF-IDF ($\uparrow$)
 Human Review 7.69
@@
 OLMo-7B (Groeneveld et al., 2024) 43.25 40.38 47.04 2.45
-Falcon-40B (Almazrouei et al., 2023) 27.34 25.13 30.88 1.06
-Gemma 2-27B (Gemma Team, 2024) 35.85 34.68 37.91 1.43
 Mistral-7B (Jiang et al., 2023) 42.03 43.80 40.77 1.17
@@
 Claude 3.5 sonnet (Anthropic, 2024a) 47.85 41.97 56.00 3.91
-GPT-4 (OpenAI et al., 2023) 47.66 42.15 55.19 5.31
 GPT-4o (OpenAI, 2024a) 47.73 42.09 55.48 5.95
@@
 AI-SCI (GPT-4o) (Lu et al., 2024) 45.05 40.02 51.91 2.23
+$\mathcal{Q}_{4}$: Do more contexts boost performance?
+We also investigate the impact of input context length for ExpDesign.
+As shown in Figure 5, we scale up the input pre-experiment context length from 0.1k to 10k tokens (10k is the length of the longest paper).
+For the experiment design, more input context does improve the performance of different LLMs, while this benefit stops after exceeding 8k tokens, which means that after the necessary information has been covered, scaling context becomes inefficient.
+Meanwhile, the explanation generation results reveal that LLMs primarily depend on given experiments rather than paper context to explain motivations.
+However, we do not expect this as we hope LLMs can explain the motivation based on a thorough understanding of the paper, just like how human experts do.
+Hence, there is still a considerable gap between the LLMs and humans in terms of grasping research motivations.
+$\mathcal{Q}_{5}$: Does multi-modal input boost performance?
+Intuitively, besides the text, when designing experiments for a given research topic, the figures can provide rich supplementary information, such as an algorithm illustration that can help better understand this research topic and underlying motivations.
+Hence, we test the performance of different LMMs (Large Multimodal Models), including GPT4-o and InternVL2 (Chen et al., 2024b).
+Table 14 shows the ablation results on the figure data.
+To our surprise, the figure data doesn’t improve the LMMs’ results in this task, even harming the performances.
+This might be due to the low informativeness of the figures, as figures usually consume more input tokens but act only as supplementary information to the text, indicating future work on developing LMMs that can effectively leverage the scientific figures.
+5.3 PaperWeakness
+Settings.
+Intuitively, full paper content is necessary for paper reviewing.
+Therefore, instead of setting a maximum input length, in Weakness, we try to utilize the whole paper.
+As the input length of Weakness is extremely long (see Table 11), we adopt a “split-combine” method — we first split the whole paper into smaller pieces and let LLMs predict the weaknesses of each piece separately; after that, we merge all pieces’ weaknesses as a final prediction.
+For the length of each small piece, we set 2,000 and 3,000 words for open- and closed-source LLMs, respectively.
+Additionally, in this task, we also examine the performance of AI-SCI (Lu et al., 2024), which enhances LLMs’ paper review ability by leveraging advanced prompting techniques, e.g., self-reflection (Shinn et al., 2024) and response ensembling (Wang et al., 2023).
+[^6] We don’t run AI-SCI on ExpDesign, because AI-SCI takes model-generated ideas as the inputs, which are incompatible with our task setting.
+Main results.
 Table 6 shows the main results, where the closed-source LLMs’ overall performances are generally superior to the results of open-source LLMs.
-Similarly, closed-source LLMs are particularly excellent in SN-Recall because of more generated weaknesses.
+Similarly, closed-source LLMs are particularly excellent in S-Recall because of more generated weaknesses.
 However, there is still a considerable gap in the weakness diversity between the LLMs and human experts.
-[^5] Note that the human’s ITF-IDF score in Table 6 can be slightly underestimated.
-This is because we keep the repeated weaknesses in the human review, which affects the human review’s informativeness (lower ITF) but is useful when calculating the SN-Recall for LLMs.
+[^7] The human’s ITF-IDF score in Table 6 can be slightly underestimated.
+This is because we keep the repeated weaknesses in the human review, which affects the human review’s informativeness (lower ITF) but is useful when calculating the S-Recall for LLMs.
 Compared with human review, most LLM-generated weaknesses are vague and lack the necessary knowledge about some frontier research works.
 Surprisingly, AI-SCI performs worse than backbone GPT-4o, especially on ITF-IDF, which suggests the challenge of Weakness, i.e., simply adopting popular prompting techniques cannot well address this task.
-Table 7: The performance comparison of different input processing methods for Weakness.
-We use GPT-4o and GPT-4-Turbo because both accept a maximum of 128k tokens input.
-We also put the results of AI-SCI in the table for reference.
-Here, “split-combine” splits the input paper into several pieces, where each piece’s length is denoted as “window size”; “no-split” means the conventional input cutting, for example, if the window size is 3,000, then only the first 3,000 words in the paper are used.
-According to the data statistics, 20,000 words can cover maximum lengths of more than 95% of the papers in our dataset.
-Models Input Context Processing Window Size (in words) SN-F1 SN-Precision SN-Recall ITF-IDF
-GPT-4-Turbo split-combine 3,000 47.66 42.15 55.19 5.31
-no-split 3,000 45.80 43.66 48.39 5.58
-no-split 20,000 44.99 42.64 47.82 5.58
-GPT-4o split-combine 3,000 47.73 42.09 55.48 5.95
-no-split 3,000 45.74 43.45 48.54 5.92
-no-split 20,000 45.47 42.97 48.51 6.02
-AI-SCI split-combine 3,000 45.05 40.02 51.91 2.23
-no-split 3,000 42.56 40.90 44.65 2.53
-no-split 20,000 42.53 40.75 44.78 2.58
 $\mathcal{Q}_{1}$: Is the split-combine effective?
-Ideally, if the LLM has a sufficient context window size, it is unnecessary to split the input papers for separate processing.
+Ideally, if the LLM has a sufficient context window size, splitting the input papers for separate processing is unnecessary.
 Consequently, in this paragraph, we utilize the LLMs accepting long context input to compare “split-combine” with “no-split”, i.e., letting LLMs write weaknesses by giving the full paper.
@@
 $\mathcal{Q}_{2}$: Does multi-modal input boost performance?
-Table 8: The ablation study about the paper tables and figures of Weakness.
-Based on the conclusion in Table 7, we use the “split-combine” to process the text input here (2,000 and 3,000 words context window size for open- and closed-source models).
-For GPT-4o, we use all the table/figure images; while for InternVL2, we randomly select two images per paper, i.e., two random figures, two random tables, or one random figure + table.
-Models SN-F1 SN-Precision SN-Recall ITF-IDF
-GPT-4o 47.73 42.09 55.48 5.95
-w/ tables 46.76 41.32 54.17 5.53
-w/ figures 46.62 41.20 54.04 5.48
-w/ tables & figures 46.58 41.17 53.98 5.36
-InternVL2-26B 41.91 41.02 43.28 1.48
-w/ tables 40.55 40.37 42.91 1.46
-w/ figures 42.88 42.10 43.76 1.46
-w/ tables & figures 42.44 42.00 43.31 1.44
 Our dataset covers both tables and figure illustrations extracted from the paper PDF as inputs.
@@
 6 Conclusion
-In this work, we propose AAAR-1.0, a novel benchmark targeting a comprehensive evaluation of the current LLMs’ AI research capacity.
-AAAR-1.0 consists of distinct expertise-intensive tasks along with the curated evaluation metrics, and collect high-quality data by employing senior AI researchers.
+In this work, we propose AAAR-1.0, a novel benchmark targeting a comprehensive evaluation of the current LLMs’ capacity in assisting AI research.
+AAAR-1.0 consists of distinct expertise-intensive tasks along with the curated evaluation metrics.
+We collect high-quality data by employing senior AI researchers and conducting strict data examinations.
 Extensive experiments highlight the challenges and values of AAAR-1.0.
-Limitations
-We shed light on two limitations of this work: i) As we gather data from open-source platforms such as arXiv and OpenReview, there is a possibility that current or future LLMs may be trained on the same source data utilized in our benchmark.
-This situation could influence the fairness of LLM comparisons and the conclusions drawn from this paper.
-While we acknowledge this potential data leakage, we maintain that our work can provide valuable insights and serve as an upper bound for some LLMs, particularly if they are indeed pretrained on those scientific papers.
-At present, this research can inform the design of current LLMs and agents, enhancing the community’s understanding of the strengths and limitations of using LLMs in scientific research.
-ii) Meanwhile, due to the high consumption of employing senior researchers in conducting data annotation, the data size for some tasks, such as ExpDesign, is relatively small.
-This might introduce variability in LLM performance, even with repeated runs.
-As this work marks the beginning of a series, we plan to release larger-scale datasets that will cover more diverse research tasks in AAAR-2.0.
-Ethics Statement
-Our study highlights that LLMs can be used to assist humans in AI research.
-This doesn’t mean we encourage the use of AI to replace human researchers.
-In contrast, we emphasize that the main responsibility for conducting scientific research should always lie with humans to prevent any societal risks, while the LLMs are only tools for making human research more efficient.
-To this end, we commit to careful distribution of the data collected in our research, ensuring it serves strictly for research purposes.
-Our goal is to mitigate the risks while maximizing the benefits offered by LLMs.
-Acknowledgments
-We would like to thank arXiv and OpenReview for releasing the paper source packages and the review comments.
-The authors would also like to thank Ibraheem Moosa and Sarkar Snigdha Sarathi Das for assisting in the data collection.
+Impact Statement
+Our study explores whether LLMs can assist human researchers in AI research.
+We do not advocate for AI replacing human researchers.
+Instead, we stress that the primary responsibility for scientific research should remain with humans to prevent societal risks, with LLMs serving as tools to enhance research efficiency.
+Specifically, our work analyzes the strengths and weaknesses of LLMs to ensure researchers remain judicious in their use of these tools.
+Our goal is to mitigate risks while maximizing the benefits offered by LLMs.
+We are committed to the careful distribution of data collected in our research, ensuring it is used solely for research purposes.
 Appendices
@@
 Appendix A Data Statistics and Diversity
-We provide the detailed data statistics of three datasets in our benchmark, as shown in Table 11, 12, and 13.
+We provide the detailed data statistics of three datasets in our benchmark, as shown in Table 9, 10, and 11.
 We use the NLTK package [^9] https://www.nltk.org/ to tokenize words and count the length.
@@
 […]
-# of classification instances 1,049
+# of positive equations 1,049
+# of negative equations 3,147
 # of source papers 869
@@
 Appendix C More Experiment Results and Details
-C.1 Human Evaluation on LLM-Generated Explanation
+C.1 Input Context Scaling Investigation
+Figure 4, Figure 5, and Table 12 show the context scaling results of EqInfer, ExpDesign, and Weakness.
+Table 12: The performance comparison of different input processing methods for Weakness.
+We use GPT-4o and GPT-4-Turbo because both accept a maximum of 128k tokens input.
+We also put the results of AI-SCI in the table for reference.
+Here, “split-combine” splits the input paper into several pieces, where each piece’s length is denoted as “window size”; “no-split” means the conventional input cutting, for example, if the window size is 3,000, then only the first 3,000 words in the paper are used.
+According to the data statistics, 20,000 words can cover maximum lengths of more than 95% of the papers in our dataset.
+Models Input Context Processing Window Size (in words) S-F1 S-Precision S-Recall ITF-IDF
+GPT-4o split-combine 3,000 47.73 42.09 55.48 5.95
+no-split 3,000 45.74 43.45 48.54 5.92
+no-split 20,000 45.47 42.97 48.51 6.02
+AI-SCI split-combine 3,000 45.05 40.02 51.91 2.23
+no-split 3,000 42.56 40.90 44.65 2.53
+no-split 20,000 42.53 40.75 44.78 2.58
+Figure 4: The input context length scaling trend on the EqInfer task.
+Figure 5: The input context length scaling trend of different LLMs on the ExpDesign task.
+C.2 Human Evaluation on LLM-Generated Novel Experiments
+Figure 6: The human guideline for evaluating the LLM-generated novel experiments.
+Figure 6 illustrates the evaluation guideline for novel experiments generated by LLMs.
+We ask 3 senior PhD students to evaluate each paper; that is, if the first two annotators disagree with each other, a third annotator will make a final decision.
+Table 13 presents several human evaluation cases.
+Table 13: Examples of human evaluation on the model-generated novel experiments.
+Paper Title Original Experiments (by human) Novel Experiment (by LLMs) Rating
+WiCE: Real-World Entailment for Claims in Wikipedia 1.
+Analysis in Verification Problem Distribution: This paper should provide detailed analysis and statistics about the verification problems in the proposed dataset.
+2.
+Off-the-shelf entailment classification performance: The authors should provide entailment classification performance of existing models on the proposed dataset without fine-tuning.
+3.
+Human Performance: The authors should show human performance on the proposed dataset.
+4.
+Performance of fine-tuned models: The authors should provide the performance of models fine-tuned on the proposed dataset.
+5.
+Performance on the evidence retrieval task: The authors should show the performance on the evidence retrieval task, which is a sub-task of the proposed dataset.
+6.
+Performance of LLMs: The authors should provide the performance of LLMs on the proposed dataset.
+7.
+Retrieval+Entailment: Authors should provide experiments on a framework of retrieving evidence sentences and evaluate entailment by using the retrieved sentences.
+8.
+Analysis of Claim-Split on Downstream Tasks: The authors should analyze how claim-split, the proposed method, is effective on tasks other than the proposed dataset.
+Assess model performance on WiCE without fine-tuning to test domain generalization from traditional NLI datasets.
+A
+MetaMath: Bootstrap Your Own Mathematical Questions for Large Language Models 1.
+Results of multiple LLMs on popular math datasets: The authors should show the performance of multiple LLMs fine-tuned on their dataset on popular math datasets.
+2.
+Performance on open-source models with different sizes: The authors should show the performance of models with different sizes trained on the proposed dataset.
+3.
+Comparison to SOTA closed-source models: The authors show compare the performance of open-source models trained on the proposed dataset and strong close-source models.
+4.
+Evaluate the effect of augmentations: The authors need to perform an ablation study to compare the different argumentation methods they proposed.
+5.
+Analyze Training on Incorrect Answers: The authors should analyze whether wrong answers generated in data augmentation can harm the performance.
+6.
+Evaluate other ways to increase the size of training data: The authors should evaluate other ways to increase the training data size and compare the performance with models trained on their proposed train data.
+7.
+Error Analysis: The authors should analyze the performance of their models in different conditions (e.g., lengths of questions).
+Prompt Sensitivity Analysis: Evaluate the sensitivity of MetaMath to different prompt formats or phrasings of mathematical questions.
+B
+Large Language Models Cannot Self-Correct Reasoning Yet 1.
+Self-Correction with Oracle Labels: The authors should evaluate self-correction performance with oracle labels.
+2.
+Intrinsic Self-Correction: The authors should show performance without using the oracle labels.
+3.
+Analysis of Mistakes in Self-Correction: The authors should analyze the properties of mistakes made in the self-correction framework.
+4.
+Multi-Agent Debate: The authors should evaluate self-correction with multi-agent debate.
+5.
+Prompt Design Analysis: The authors should analyze the influence of prompt design for the initial responses on self-correction performance.
+Visualization of learned representations or attention mechanisms to provide insights into the model’s inner workings.
+C
+C.3 Human Evaluation on LLM-Generated Explanation
 We ask 5 annotators to evaluate the LLM-generated explanations.
@@
 For each paper, the annotator is given 5 different models’ outputs (model names are anonymized), and the annotator has to decide if each LLM-generated explanation is acceptable according to the experiment.
-We show the human evaluation results in Table 4,
+We show the human evaluation results in Table 5.
+C.4 Multi-Modal Input Ablation
+We post the multi-modal ablation study of ExpDesign and Weakness in Table 14 and Table 15.
+Table 14: The figure inputs ablation of ExpDesign.
+For the maximum text input length, same as the setting in Table 2, we use 2,000 and 3,000 words for open- and closed-source models, respectively.
+For the closed-source GPT-4o and GPT-4, as they have long context window sizes, we use all the figures of each paper.
+While for InternVL2, we randomly select two figures per input paper.
+Experiment Design Experiment Explanation
+Models En-F1 En-Precision En-Recall S-Match ROUGE-L ROUGE-1
+GPT-4o 25.03 22.25 36.59 58.54 29.25 35.50
+w/ figures 25.39 24.35 32.80 58.53 27.87 34.30
+InternVL2-26B 24.26 39.50 14.91 50.03 29.13 34.26
+w/ figures 15.04 38.50 8.64 50.29 29.29 34.06
+Table 15: The ablation study about the paper tables and figures of Weakness.
+Based on the conclusion in Table 12, we use the “split-combine” to process the text input here (2,000 and 3,000 words context window size for open- and closed-source models).
+For GPT-4o, we use all the table/figure images; while for InternVL2, we randomly select two images per paper, i.e., two random figures, two random tables, or one random figure + table.
+Models S-F1 S-Precision S-Recall ITF-IDF
+GPT-4o 47.73 42.09 55.48 5.95
+w/ tables 46.76 41.32 54.17 5.53
+w/ figures 46.62 41.20 54.04 5.48
+w/ tables & figures 46.58 41.17 53.98 5.36
+InternVL2-26B 41.91 41.02 43.28 1.48
+w/ tables 40.55 40.37 42.91 1.46
+w/ figures 42.88 42.10 43.76 1.46
+w/ tables & figures 42.44 42.00 43.31 1.44
 Appendix D Data cases and Annotation Platform Illustration
-As shown in Figure 7, 8, and 9, we show the sample cases of the three tasks in AAAR-1.0. Meanwhile, we illustrate the screenshot of our annotation platform in Figure 6.
+As shown in Figure 8, 9, and 10, we show the sample cases of the three tasks in AAAR-1.0. Meanwhile, we illustrate the screenshot of our annotation platform in Figure 7.
 Figure 7: The annotation platform for collecting the annotation of ExpDesign.
@@
 Appendix E Prompt Templates
-In this appendix, we attach all the prompts used in this work, including prompts in data collection and model prediction, as shown in Figure 10, 11, and 12.
-Figure 7: A sample case of EqInfer.
+In this appendix, we attach all the prompts used in this work, including prompts in data collection and model prediction, as shown in Figure 11, 12, and 13.
+Figure 8: Two sample cases of EqInfer.
 Figure 9: A sample case of ExpDesign.

### v4

@@
 However, researchers face unique challenges and opportunities in leveraging LLMs for their own work, such as brainstorming research ideas, designing experiments, and writing or reviewing papers.
-In this study, we introduce AAAR-1.0, a benchmark dataset designed to evaluate LLM performance in three fundamental, expertise-intensive research tasks: (i) EquationInference, assessing the correctness of equations based on the contextual information in paper submissions; (ii) ExperimentDesign, designing experiments to validate research ideas and solutions; (iii) PaperWeakness, identifying weaknesses in paper submissions; and (iv) ReviewCritique, identifying each segment in human reviews is deficient or not.
+In this study, we introduce AAAR-1.0, a benchmark dataset designed to evaluate LLM performance in four fundamental, expertise-intensive research tasks: (i) EquationInference, assessing the correctness of equations based on the contextual information in paper submissions; (ii) ExperimentDesign, designing experiments to validate research ideas and solutions; (iii) PaperWeakness, identifying weaknesses in paper submissions; and (iv) ReviewCritique, identifying each segment in human reviews is deficient or not.
 AAAR-1.0 differs from prior benchmarks in two key ways: first, it is explicitly research-oriented, with tasks requiring deep domain expertise; second, it is researcher-oriented, mirroring the primary activities that researchers engage in on a daily basis.
@@
 To this end, in this work, we introduce AAAR-1.0, a novel benchmark that aims to comprehensively assess the LLMs’ capacity on expert-level research tasks.
-As illustrated in Figure 1, AAAR-1.0 decomposes three distinct expert-level AI research tasks from the researcher’s daily activities, including i) EquationInference, investigating whether the LLMs can infer the equation correctness based on the paper context; ii) ExperimentDesign, validating LLMs’ ability on designing reliable experiments for a research idea; iii) PaperWeakness, testing the quality of weaknesses discovered by LLMs from paper drafts; and iv) ReviewCritique, investigating whether LLMs can identify and explain the deficient/unreliable human-written paper reviews.
+As illustrated in Figure 1, AAAR-1.0 decomposes four distinct expert-level AI research tasks from the researcher’s daily activities, including i) EquationInference, investigating whether the LLMs can infer the equation correctness based on the paper context; ii) ExperimentDesign, validating LLMs’ ability on designing reliable experiments for a research idea; iii) PaperWeakness, testing the quality of weaknesses discovered by LLMs from paper drafts; and iv) ReviewCritique, investigating whether LLMs can identify and explain the deficient/unreliable human-written paper reviews.
 To ensure data quality, senior AI researchers with extensive domain expertise perform data annotation for AAAR-1.0, followed by rigorous multi-round data examination and filtering.
-All three tasks require models to possess strong domain knowledge covering various cutting-edge research findings, as well as expert-level research experience, to the extent that even humans need substantial research accumulation to tackle the tasks we designed.
+All tasks require models to possess strong domain knowledge covering various cutting-edge research findings, as well as expert-level research experience, to the extent that even humans need substantial research accumulation to tackle the tasks we designed.
 Crucially, tasks here are singular, stand-alone challenges (with clear input and output expectations) rather than a complicated task chain (Li et al., 2024; Lu et al., 2024), providing a more transparent assessment of the model’s intermediate output.
@@
 LLMs for AI Research.
-With the rapid evolution of pertaining techniques, LLMs are found to be useful in assisting various research disciplines (Yu et al., 2024a; Labrak et al., 2024), particularly in AI research, such as generating novel research ideas (Kumar et al., 2024; Yu et al., 2024b), reviewing research draft (Gao et al., 2024; Du et al., 2024; Liang et al., 2024), and writing scientific papers (Chamoun et al., 2024; Lu et al., 2024).
+With the rapid evolution of pertaining techniques, LLMs are found to be useful in assisting various research disciplines (Yu et al., 2024a; Labrak et al., 2024), particularly in AI research, such as generating novel research ideas (Kumar et al., 2024; Yu et al., 2024b), reviewing research draft (Gao et al., 2024; Du et al., 2024; Liang et al., 2024; Zhu et al., 2025), and writing scientific papers (Chamoun et al., 2024; Lu et al., 2024; Weng et al., 2024).
 For example, Si et al.
@@
 However, there is still a lack of systematic evaluations and quantitative analyses on the LLMs’ (intermediate) output of each single-step research task.
-Our work focuses on building a benchmark that has individual research steps with clear input-output expectations, thus making it suitable for comprehensive LLMs evaluation.
+Accordingly, our work focuses on building a benchmark consisting of individual research steps with clear input-output expectations, making it suitable for comprehensive LLM evaluation.
+Moreover, we emphasize that relying on LLMs to fully replace human effort might compromise academic integrity.
+While our benchmark primarily serves an educational purpose — LLMs assist junior researchers by providing imperfect but insightful ideas, rather than by governing the entire research process.
 Benchmarks for AI Research Tasks.
@@
 o1-preview (OpenAI, 2024b) 46.35 31.43 88.27
+o3-mini (OpenAI, 2025) 47.98 34.34 79.59
 Settings.
@@
 Notably, though the performance of Mixtral is slightly superior to the baseline, the extremely biased precision and recall scores imply that Mixtral is also simply predicting almost all samples as positive instead of truly inferring.
-Meanwhile, compared to the All-Positive baseline, the performance superiority of the strong close-source LLMs is not significant, the best LLM on this task only obtains 46.74%, which demonstrates the challenge of EqInfer compared with other similar benchmarks (Song et al., 2023).
+Meanwhile, compared to the All-Positive baseline, the performance superiority of the strong close-source LLMs is not significant, the best LLM on this task only obtains 47.98%, which demonstrates the challenge of EqInfer compared with other similar benchmarks (Song et al., 2023).
 The generally high recall with low precision of all LLMs also indicates real-world risks, e.g., relying on LLMs to check the validity of equations in paper review.
@@
 o1-preview (OpenAI, 2024b) 30.13 28.13 38.59 58.55 29.11 36.70
+o3-mini (OpenAI, 2025) 30.17 28.70 37.67 54.01 20.71 29.14
 Settings.
@@
 o1-preview (OpenAI, 2024b) 48.62 42.54 57.08 5.63
+o3-mini (OpenAI, 2025) 46.33 42.00 51.99 5.85
 LLM Agent Framework
@@
 Extensive experiments highlight the challenges and values of AAAR-1.0.
+Limitations
+We shed light on two limitations of this work: i) As we gather data from open-source platforms such as arXiv and OpenReview, there is a possibility that current or future LLMs may be trained on the same source data utilized in our benchmark.
+This situation could influence the fairness of LLM comparisons and the conclusions drawn from this paper.
+While we acknowledge this potential data leakage, we maintain that our work can provide valuable insights and serve as an upper bound for some LLMs, particularly if they are indeed pretrained on those scientific papers.
+At present, this research can inform the design of current LLMs and agents, enhancing the community’s understanding of the strengths and limitations of using LLMs in scientific research.
+ii) Meanwhile, due to the high consumption of employing senior researchers in conducting data annotation, the data size for some tasks, such as ExpDesign, is relatively small.
+This might introduce variability in LLM performance, even with repeated runs.
+As this work marks the beginning of a series, we plan to release larger-scale datasets that will cover more diverse research tasks in AAAR-2.0.
+Acknowledgments
+We would like to thank arXiv and OpenReview for releasing the paper source packages and the review comments.
+The authors would also like to thank Ibraheem Moosa and Sarkar Snigdha Sarathi Das for assisting in the data collection.
 Impact Statement